### v1

# Funny Valen-Tine: Solving visual abstract reasoning problems through defining the solution distribution

Abstract
The importance of visual abstruct reasoning problems in the field of image processing is beyond words. Both Bongard-Logo problems and Raven’s Progressive Matrices (RPM) belong to the domain of visual abstract reasoning tasks, with Bongard-Logo categorized as image clustering reasoning and RPM involving image progression pattern reasoning. This paper introduces a novel baseline model, Valen, which falls under the umbrella of probabilistic highlighting models. Valen demonstrates remarkable performance in solving both RPM and Bongard-Logo problems, offering a versatile solution for these reasoning tasks. Our investigation extends beyond the application of Valen, delving into the underlying mechanisms of probability-highlighting solvers when tackling visual abstract reasoning problems. We revisit how these solvers handle instances in tasks like RPM and Bongard-Logo, realizing that they approximate the solution to each reasoning problem instance as a distribution delineated by primary and auxiliary samples. This led us to propose that the learning objective of probability-highlighting solvers is not the distribution of correct solutions but rather a distribution defined by both primary and auxiliary samples. To bridge the discrepancies, we introduced the Tine method, an adversarial learning-based approach designed to assist Valen in estimating a solution distribution closer to the correct one. However, Tine’s adversarial framework also introduces issues such as unstable training. Reflecting on Tine, we propose modeling the sample distribution of reasoning problems as a mixture of Gaussian distributions, leading to the development of the Funny method. Funny effectively models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution. Furthermore, we designed the SBR method to model the distribution of progressive patterns representation as a mixture of Gaussian distributions. Overall, the Funny, Tine, and SBR methods have significantly improved Valen’s performance, providing new ideas and methods for the study of visual abstract reasoning problems.
Index Terms:
Abstract reasoning, Raven’s Progressive Matrices, Bongard-Logo problem, Adversarial Learning, Gaussian Mixture Model.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Deep neural networks have made remarkable achievements in various fields, including computer vision[1, 2, 3], natural language processing[4, 5, 6], generative model[7, 8, 9], visual question answering[10, 11], and visual abstract reasoning[14, 12, 13].
Visual abstract reasoning plays a pivotal role in the field of image processing. It is not only one of the advanced tasks in image processing, requiring systems to recognize objects in images and understand the relationships between them for logical inference, but also a key driver for technological innovation and application expansion in image processing. By addressing visual abstract reasoning problems, new algorithms and techniques such as deep learning and graph convolutional networks can be explored and applied, thereby propelling extensive applications and developments in fields like autonomous driving, medical diagnosis, and security surveillance. However, this field still faces challenges such as insufficient and imbalanced data, as well as limited model interpretability. Future research will focus more on enhancing the robustness, real-time performance, and interpretability of algorithms to further elevate the potential and value of visual abstract reasoning technology.
The RPM problem [14] and the Bongard problem [12, 13] are two renowned examples within the realm of visual abstract reasoning. These two problems are the target problems that this paper aims to solve.
I-A RAVEN and PGM
RPM[14], or Raven’s Progressive Matrices, is a classic visual reasoning intelligence test typically presented as a 3x3 matrix. It requires test-takers to observe and analyze visual features such as shape, size, direction, position, and their relationships within the problem statement, identify patterns or rules, and select an answer from given option pool that completes the matrix. Due to its non-verbal nature and value in assessing general intelligence, RPM is widely used in psychology, education, neuroscience, and other fields, serving as a tool for evaluating intelligence levels and studying human cognitive processes, intellectual development, and brain function. With the advancement of AI technology, researchers are exploring computer algorithms to solve RPM problems, albeit facing challenges.
RAVEN[15] and PGM[16] are classic examples of RPM problems, with Figure 1 illustrating these two types of problems respectively.
Figure 1: RAVEN instance (a) and PGM instance (b)
I-B Bongrad-Logo
Distinct from the image progressive pattern reasoning problems like RPM, the Bongard problem [12, 13] belongs to the category of image clustering reasoning problems. It typically involves multiple images divided into primary and auxiliary groups, with images in the primary group sharing an abstract concept governed by a specific rule and those in the auxiliary group deviating from these rules. Solving Bongard problems requires deep learning algorithms to accurately classify ungrouped images. As a specific instance within abstract reasoning, the Bongard-Logo problem [13] poses a significant challenge due to its inferential complexity. Each case comprises 14 images: six from the primary group, six from the auxiliary group, and two for classification. The grouping criterion relies on geometric shapes and their arrangement. Figure 2 illustrates a Bongard-Logo case.
Figure 2: Bongard-Logo instance
II Related work
II-A RPM solvers and Bongard-Logo solvers
There are numerous excellent solvers for RPM (Raven’s Progressive Matrices) and Bongard-Logo problems, among which CoPINet[19], LEN+teacher[18], DCNet[20], NCD[22], SCL[23], SAVIR-T[24], PrAE[25], NVSA[27], ALANS[26], RS-TRAN[28] and CRAB[29] are notable RPM solvers, while SBPDL[30], PMoC[31] and Triple-CFN[32] are outstanding Bongard-Logo solvers. Since the method proposed in this paper is relatively independent and does not have a strong dependency on these existing works, the details about them will not be elaborated further.
III Methodology
Both Bongard-Logo problems and Raven’s Progressive Matrices (RPM) fall under the umbrella of visual abstract reasoning tasks. Specifically, Bongard-Logo problems can be categorized as image clustering reasoning, while RPM challenges involve image progression pattern reasoning.
This paper argues that explicitly defining the distribution of solutions for the designed abstract reasoning problem solver is a straightforward and effective approach. To this end, we first designed a general solver, Valen, for both RPM and Bongard-Logo problems. Additionally, we developed three accompanying methods: Funny, Tine, and SBR. These three methods all aim to explicitly define the distribution of correct solutions for Valen through explicit or implicit approaches.
Here we clarify a definition: in this paper, samples that conform to the underlying patterns within instances of reasoning problems are referred to as “primary samples.” For example, the correct options within RPM problem instances and the samples in the primary group within Bongard-Logo problem instances are considered primary samples. Conversely, samples that do not conform to the underlying patterns, such as the incorrect options in RPM problem instances and the samples in the auxiliary group within Bongard-Logo problem instances, are referred to as “auxiliary samples.”
IV Valen: a new baseline model for RPM and Bongard-Logo problems
In this section, we introduce Valen (Visual Abstraction Learning Network), our baseline model designed to address both RPM (Raven’s Progressive Matrices), an image progressive pattern reasoning problem, and Bongard-Logo, an image clustering reasoning problem.
IV-A Valen targeting RPM problem
The RPM problem challenges participants to complete a 3×3 matrix of images by choosing answers from a pool of options. A correct answer is one that, when placed in the matrix, maintains a consistent and specific image progression pattern that can be deduced from the incomplete matrix, referred to as the problem statement. [39].
This paper requires a baseline to support our reasoning work. This baseline is expected to determine the plausibility of an RPM matrix based on the progression patterns exhibited within it and assign a probability to reflect this plausibility.
In RPM problems, there exists a unique design where an incomplete RPM matrix (missing one image) can still convey a complete progression pattern. This design is essential as RPM serves as an assessment tool. Hence, this paper asserts that the reasoning process of a successful RPM problem solver should be schematically represented as illustrated in the Figure 3.
Figure 3: The reasoning process of a envisioned RPM problem solver
As seen in the figure, our envisioned RPM solver consists of two modules: the Progressive Pattern Extractor and the Pattern Consistency Analyzer. Additionally, our envisioned RPM solver can perform the following three reasoning processes: 1. The process indicated by the blue line: When the RPM problem statement is completed by an option from the option pool to form a full $3\times 3$ progression matrix, the solver can enumerate all $N$ types of incomplete matrix of this $3\times 3$ matrix. 2. The process denoted by the green line: For each enumerated incomplete matrix, the Progressive Pattern Extractor extracts the progression pattern, resulting in a total of $N$ sets of progression patterns. 3. The process marked by the orange line: The solver utilizes its Pattern Consistency Analyzer to evaluate the consistency of these $N$ sets of progression patterns and outputs a score to quantify this consistency. This score also serves as the score for the corresponding option. The aforementioned three processes will also constitute the procedure of Valen.
Since this paper provides a clear definition of the reasoning process for a successful RPM solver, our designed Valen is derived by refining these three processes. Specifically, our Valen also consists of three processes: representation extraction of matrix and enumeration of incomplete matrix, extraction of progressive patterns from incomplete matrix, and consistency analysis of progressive patterns, which correspond to the aforementioned three processes respectively. We will introduce these three refined processes one by one.
IV-A1 Representation extraction of matrix and enumeration of incomplete matrix
In this paper, this process in Valen is designed as the structure shown in Figure 4.
Figure 4: Representation extraction of matrix
As shown in figure 4, when the problem statement of the RPM task is completed by a specific option to form a $3\times 3$ progression matrix, denoted as $\{x_{i}|i\in[1,9]\}$, Valen utilizes a basic Vision Transformer (ViT) as a feature extractor to obtain multi-viewpoint representations of $\{x_{i}|i\in[1,9]\}$. These representations are denoted as $\{r_{i}|i\in[1,9]\}_{v}$, where $v\in[1,V]$ and $V$ represents the number of viewpoints, $v$ represents the index of the viewpoint, and $i$ represents the index of the image in the matrix. The subsequent processes in Valen will handle these multi-viewpoint representation matrices in parallel and equally. The method of processing RPM images into multi-viewpoint representations has been applied in previous works [28, 24].
Subsequently, these multi-viewpoint representations $\{r_{i}|i\in[1,9]\}_{v}|_{v=1}^{V}$ will be enumerated by Valen to generate all possible incomplete representations matrix. The enumeration results are illustrated in the Figure 5, with the last item of the enumeration is the matrix with a incomplete option position.
Figure 5: Enumeration results of incomplete matrix
We denote these enumerated multi-viewpoint incomplete representation matrices as $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$, where $n$ represents the index of the incomplete representation matrix, and $i$ still represents the number of the representation within the incomplete matrix. Please note the order of the enumeration results set in the figure.
IV-A2 Extraction of progressive patterns from incomplete matrix
In Valen, this process is undertaken by a Transformer-Encoder. Specifically, we use a standard Transformer-Encoder to process all the previously enumerated multi-viewpoint incomplete representation matrices, with the processing procedure illustrated in Figure 6.
Figure 6: Extraction of progressive patterns from incomplete matrix
The figure illustrates the process of the Transformer-Encoder handling $\{r_{9i}|i\in[1,8]\}_{v}$, which represents the last item of the enumeration results of incomplete representation matrices from a specific viewpoint $v$. The process of handling other incomplete representation matrices $\{r_{ni}|n\in[1,8]i\in[1,9],i\neq n\}_{v}$ is similar to the one shown in the figure, with no essential differences.
In the figure 6, we can observe that we use an MLP to extract all row and column progressive information from the incomplete representation matrix $\{r_{9i}|i\in[1,8]\}_{v}$. Subsequently, we process all four row and column progressive information using a standard Transformer-Encoder to obtain $M$ representations of progressive patterns. $M$ is a hyperparameter that can be freely set, and in this paper, it is set to 2. The impact of setting $M$ to other values on Valen is not explored in this paper. This asymmetric seq-to-seq implementation (four-to-$M$) has been introduced in related techniques of ViT[17]. After processing all the incomplete representation matrices $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$ using the standard Transformer-Encoder, we obtain all multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$.
IV-A3 Consistency analysis of progressive patterns
In this process, Valen will evaluate the consistency among these multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$. The task of this consistency analysis will be accomplished by the Transformer-Decoder. The process of evaluating the consistency of the progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$ from a specific viewpoint $v$ can be conducted using a Transformer-Decoder, as illustrated in Figure 7.
Figure 7: Evaluating the consistency of the progressive patterns representation $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$
In the figure 7, we can observe that the progressive pattern representations $\{P_{9m}|m\in[1,M]\}_{v}$ extracted from the incomplete representation matrix $\{r_{9i}|i\in[1,8]\}_{v}$ are used as queries, while the other progressive pattern representations $\{P_{nm}|n\in[1,8],m\in[1,M]\}_{v}$, which are extracted from the incomplete representation matrices $\{r_{ni}|n\in[1,8],i\in[1,9],i\neq n\}_{v}$, serve as key-value pairs input into the Transformer-Decoder to compute the output vector set. The $M$ vectors within the output vector set are then individually mapped to a score by a new MLP. These mapped $M$ scores are summed to obtain the consistency score of the progressive patterns under viewpoint $v$.
Subsequently, We calculate the average of the progressive pattern consistency scores obtained from all viewpoints to determine the overall viewpoint score for the corresponding option. Finally, we employ cross-entropy as the loss function to constrain the scores of all options according to the logic of the RPM task, in order to optimize the learnable parameters involved in the aforementioned three processes.
IV-B Valen targeting Bongard-Logo problem
We consider Valen as a good baseline model for dealing with image progressive pattern reasoning problems, but it can also handle image clustering reasoning problems such as Bongard-Logo. To adapt Valen to the Bongard-Logo problem, we need to transform this problem accordingly.
In a particular instance of the Bongard-Logo problem, we use $x_{i}$ to represent all the images. Specifically, the images in the primary group of this instance are denoted by $\{x_{i}|i\in[1,6]\}$, while those in the auxiliary group are represented by $\{x_{i}|i\in[8,13]\}$. Furthermore, $x_{7}$ represents the test sample that should be assigned to the primary group, and $x_{14}$ represents the test sample that should be assigned to the auxiliary group.
This paper proposes that we can adopt different perspectives to view image clustering reasoning problems, such as the Bongard-Logo problem. Specifically, we can treat them as an option evaluation problem akin to the RPM problem. To elaborate, the images in the primary group $\{x_{i}|i\in[1,6]\}$ can be regarded as a problem statement similar to that in the RPM problem, while the test images and images in the auxiliary group $\{x_{i}|i\in[7,14]\}$ can be seen as an option pool anaLogous to the one in the RPM problem. This transformation process can be visualized through Figure 8.
Figure 8: Transformation of Bongard-Logo problem
After such transformation, Valen can be effectively applied to image clustering reasoning problems like the Bongard-Logo problem.
It is noteworthy that when migrating Valen to Bongard-Logo, some adjustments are still necessary to accommodate the data characteristics of Bongard-Logo. Specifically, due to the scarcity of training instances in the Bongard-Logo problem, the ViT-based image representation extractor may struggle to perform effectively[17, 28]. Furthermore, although we transform the Bongard-Logo problem into an RPM-like format, it is unnecessary to analyze the progressive patterns among Bongard-Logo images, as such patterns do not exist in clustering reasoning problems like Bongard-Logo.
Therefore, we have redesigned the image representation extractor, originally illustrated in Figure 4, into a convolutional network concatenated with a Transformer-Encoder. The new extractor design can be visualized in Figure 9.
Figure 9: Representation extraction of Bongard-Logo image
In the figure 9, we can observe that when a Bongard-Logo image is processed by the new Valen’s feature extractor, it is first encoded into a feature map by ResNet18. Subsequently, this feature map is decomposed into multiple tokens and fed into a standard Transformer-Encoder to obtain multi-viewpoint representations. Such a design allows for the combination of CNN’s ability to analyze a smaller number of samples[1, 2, 3, 17] and ViT’s ability to analyze samples from multiple viewpoint[28, 24]. This design has been applied in previous work[24].
Additionally, we have simplified the incomplete matrix progressive pattern extractor, which was originally intended to be structured as shown in Figure 6, into the form illustrated in Figure 10.
Figure 10: Pattern extraction of Bongard-Logo incomplete matrix
In the figure10, we can see that we have abandoned the use of MLP to extract the row and column progressive information from the incomplete matrix in Bongard-Logo, because the transformed Bongard-Logo problem only contains clustering patterns and not progressive patterns. Instead, we directly feed the representations of all images within the incomplete matrix of Bongard-Logo as tokens into the Transformer-Encoder to extract $M$ clustering patterns of Bongard-Logo. When Valen processes Bongard-Logo problems, $M$ is set to 6.
V Tine, a Method for Advancing the Solution Distribution Boundary
Current visual abstract reasoning problem solvers, including Valen, mostly follow a pattern: solvers are keen on assigning probabilities to each option of reasoning problems. The loss functions designed for these solvers are diverse, and the backbone network designs of these solvers vary greatly, but most of them adhere to this paradigm. We name this type of solver as “probability-highlighting solver”.
Therefore, we can interpret these solvers as a conditional probability distributions obeyed by abstract reasoning problems, or understand them as providing distributions of solutions to abstract reasoning problems. If we use RPM solvers as examples, then these solvers can be abstracted as a conditional probability distribution, denoted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$. (Accordingly, when Valen is applied to Bongard-Logo problems, it can be abstracted as a conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{6},\theta)$). In this context, the two conditions of the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ are the statement of the reasoning problem $\{x_{i}\}_{i=1}^{8}$ and the optimizable parameters $\theta$ of the solver. Here, the $x_{\alpha}$ refers to the option that is pending verification. This conditional distribution implies that, when given a problem statement $\{x_{i}\}_{i=1}^{8}$ and an option $x_{\alpha}$, it is possible to calculate the probability of $x_{\alpha}$ being a solution to $\{x_{i}\}_{i=1}^{8}$.
The optimization of probability-highlighting solver can be regarded as optimizing the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$. A wide range of loss functions can be used for optimization, but they all share the same objective: to assign high probabilities to the primary samples in the probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ and low probabilities to the auxiliary options. Such an optimization objective makes us realize that the optimization process of the probability-highlighting solver is not aimed at fitting the distribution of correct solutions to visual abstract reasoning problems, but rather is dedicated to fitting a distribution jointly delimited by primary and auxiliary samples.
To simplify the explanation of our findings, we assume that the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ follows a Gaussian distribution. Consequently, we can explicitly identify the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta})$. Then, the negative logarithm of the probability density function of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta})$ can be expressed as follows:
$-\log(P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta}))=\frac{{(x_{a}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(|\,\sigma_{\theta}|\,)$ (1)
Evidently, we have omitted the constant term $-1/2\log(2\pi)$. In this formula, $\mu_{\theta}$ and $\sigma_{\theta}$ respectively represent the mean and standard deviation of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$. Since we have defined the form of distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a Gaussian distribution, we consequently abstract the optimizable parameters $\theta$ into optimizable mean $\mu_{\theta}$ and standard deviation $\sigma_{\theta}$.
Optimizing the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta})$ using the loss function is equivalent to optimizing the values of $\mu_{\theta}$ and $\sigma_{\theta}$. Our optimization objective is that when primary samples (the correct options in an RPM problem, or the samples within the main group in a Bongard-Logo problem) are plugged into formula (1), we expect to obtain a smaller value. Conversely, when auxiliary samples (the incorrect options in an RPM problem, or the samples within the auxiliary group in a Bongard-Logo problem) are plugged into formula (1), we hope that the formula returns a larger value.
Evidently, constraining the return value of the formula to achieve a minimum is equivalent to optimizing the value of $\mu_{\theta}$ in the formula. Conversely, constraining the return value to achieve a maximum is equivalent to minimizing the value of $\sigma_{\theta}$ in the formula. In other words, the probability-highlighting solver positions the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ for the reasoning problem through primary samples, while the standard deviation of the solution distribution is constrained by auxiliary samples.
It appears to be a discouraging reality that the learning objective of a probability-highlighting solver is not to determine the distribution of the correct solutions to reasoning problems, but rather to approximate a distribution that is delimited by both primary and auxiliary samples within those problems. This suggests that the reasoning accuracy of the probability-highlighting solver is constrained. The standard deviation of the distribution delimited by the auxiliary samples is necessarily greater than or equal to the standard deviation of the correct solution distribution. In other words, the distribution delimited by both primary and auxiliary samples is often much coarser than the distribution of correct solutions. Aiming to approximate the former as the optimization objective is akin to “aiming for mediocrity and achieving even less”. This discouraging reality can be visualized in Figure 11. Note that the figure 11 is intended to schematically represent our analysis results, rather than to depict the actual form of the solution distribution or the correct solution distribution.
Figure 11: Schematic diagram of the relative relationship between the solution distribution, the correct solution distribution, and the primary and auxiliary samples.
This paper argues that to further improve the reasoning accuracy of Valen in solving RPM problems and Bongard-Logo problems, it is necessary to provide Valen with more judicious auxiliary samples. By “more judicious auxiliary samples,” we refer to those that can better delimit the solution distribution provided by Valen. These more judicious auxiliary samples have the potential to narrow down the standard deviation of the solution distribution. The visualization of these more judicious auxiliary samples can be found in Figure 12.
Figure 12: judicious auxiliary samples
As seen in the figure 12, the judicious auxiliary samples we desire are those that appear more authentic or closer to the primary samples from Valen’s perspective, compared to the original auxiliary samples provided in the reasoning problems. Since the property of being “judicious” is evaluated from Valen’s perspective, whether an auxiliary sample is judicious or not, and the extent of its judiciousness, lacks interpretability from a human perspective. This implies that, relative to the primary samples, samples with smaller attribute shifts or fewer shifted attributes may not necessarily be considered more judicious from Valen’s perspective.
Therefore, this paper argues that we can adopt an adversarial learning strategy to provide Valen with auxiliary samples that are more judicious from Valen’s perspective. Given this insight, we design Tine (Technique for edge Identification of solution distribution based on adversarial Networks) method. Specifically, we intend to emulate adversarial generation techniques similar to GAN[7] or WGAN[36] to generate more judicious samples for Valen. However, unlike GAN or WGAN, whose objective is to train robust and high-performance generators, our utilization of adversarial learning aims to refine Valen, which serves as the discriminator in this context.
Our objective is to design an auxiliary sample generator for Valen. The task of this generator is clear: to produce auxiliary samples that can achieve higher scores (probability) in Valen compared to the original auxiliary samples provided in the reasoning tasks. Conversely, Valen’s role is to classify these auxiliary samples as low-scoring ones. During the training process, the parameters of Valen and this generator are updated alternately and exclusively, with their performance gradually refined through adversarial learning, which is a common process in adversarial learning.
Given that our pursuit is to generate more judicious auxiliary samples from Valen’s perspective, we are not concerned with their interpretability, attribute geometry, or whether they strictly conform to the nature of reasoning problems. Therefore, we believe that instead of forcing the generator to produce image-based auxiliary samples, which would incur greater computational overhead and optimization difficulties, it is more effective to generate representations of judicious auxiliary samples. By solely generating multi-viewpoint representations of reasoning problems, the generator’s structure becomes simpler and easier to optimize. Consequently, our generator is designed as shown in Figure 13. The figure 13 depicts Valen, which is tailored for addressing the RPM problem, and its accompanying auxiliary sample representation generator. Since we have already transformed the Bongard-Logo problem into an RPM-like problem, we will not further elaborate on the structure of the generator for the Bongard-Logo version.
Figure 13: The structure of the generator associated with Valen
As depicted in Figure 13, the left gray dashed box encloses the previously designed Valen, while the right gray dashed box encloses the generator designed for Valen. The figure 13 shows that the backbone of this generator is a standard Transformer-Encoder, which generates representations of more judicious auxiliary samples based on the representations of the abstract reasoning problem statement $\{r_{i}|i\in[1,8]\}_{v}$, the progressive patterns $\{P_{9m}|m\in[1,M]\}_{v}$ of the statement, and $L$ learnable vectors. Both the representations of the statement and their progressive patterns are provided by Valen, while the $L$ learnable vectors require additional establishment for the generator. It is noteworthy that when dealing with PGM problems in RPM tasks, we have designed special positional encoding for the representations of the problem statement. This positional encoding can be specifically attached to incomplete matrices with missing option positions and is set symmetrically along the diagonal. After the Transformer-Encoder processes the input representations (i.e., $\{r_{i}|i\in[1,8]\}_{v}$, $\{P_{9m}|m\in[1,M]\}_{v}$ and $L$ learnable vectors) and generates the output, we extract and save those $L$ optimizable vectors from the output. These $L$ (processed) output learnable vectors are then individually processed by an MLP into vectors of double length. These $L$ doubled-length vectors are transformed into $L$ vector using the reparameterization technique[8]. The process of reparameterization, which is a practical approach used in VAE[8], can be visualized as Figure 14.
Figure 14: Reparameterization
These $L$ reparameterized vector are regarded as the representations of $L$ auxiliary samples. It is worth noting that $L$ is a hyperparameter that can be freely set, determining how many judicious auxiliary samples Tine generates for Valen in one instance. In this paper, $L$ is set to 9.
In summary, our Tine method is proposed after summarizing the working principles of probability-highlighting solvers such as Valen. We believe that such solvers establish a solution distribution for each instance of a reasoning problem, with the optimization objective being to assign high probability to primary samples and low probability to auxiliary samples. This means that the optimization target of the solution distribution is not aligned with the distribution of correct solutions to the reasoning problems, but rather with the distribution delimited by both primary and auxiliary samples. Unfortunately, this delimited distribution is often much coarser than the correct solution distribution. To make the distribution delimited by both primary and auxiliary samples closer to the correct solution distribution, we designed Tine based on adversarial learning. Although subsequent experiments demonstrated that Tine can effectively improve Valen’s performance, we believe that the Tine method still fails to address the underlying essence of the problem. Therefore, further exploration is conducted in this paper.
VI Funny: Method for Planning the Distribution of Reasoning Problem Samples.
Based on the analysis in the previous section, auxiliary samples provided in reasoning tasks constrain the standard deviation of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$, while primary samples locate the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$. The Tine method generates more judicious auxiliary samples to further constrain the standard deviation. Although experimental results demonstrate the effectiveness of Tine, it remains uncertain whether Tine constrains the standard deviation too little or not enough. Furthermore, Tine based on adversarial learning also encounters all the issues associated with adversarial learning, such as unstable training processes and difficulties in determining the completeness of training. Given this series of challenges, this paper proposes the Funny (Framework Utilizing Neural Networks for Yielding Image Representation Distribution) method.
Our Funny method argues that instead of laboriously adopting adversarial learning strategies to make the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ provided by the solver closer to the correct solution distribution, it is more effective to directly specify the range of correct solutions or, alternatively, to explicitly define the details of the correct solution distribution. This paper will next analyze the setting and determination of correct solutions for image progressive pattern reasoning problems such as RPM and image clustering reasoning problems such as Bongard-Logo.
According to the design of RPM problems, the attributes of an image within the RPM problem are divided into key attributes related to reasoning and non-key attributes unrelated to reasoning. The design purpose of attributes unrelated to reasoning is to interfere with the solver’s reasoning and test its ability to filter attributes. This paper has discovered a point that cannot be neglected: attributes are the only factor that determines the specificity of RPM images. Any two RPM images with the same key attributes have the same identity in the reasoning process, regardless of how different their non-key attributes are. This implies that any sample whose key attributes pass the verification can be considered a correct solution to the RPM problem. For instance, in the RAVEN problem[15], the angle of an entity is considered a non-key attribute that is unrelated to reasoning, while attributes such as color, size, shape, and position are all key attributes. This leads to the situation where a white, medium-sized hexagon positioned in the center retains its reasoning identity regardless of how it is rotated. No matter how it rotates, it will not be transformed from a correct solution to an incorrect one or vice versa.
The design of correct solutions for Bongard-Logo problems is much more complex. The reasoning identity of a Bongard-Logo sample is determined by the context within the specific instance it belongs to. Among the various attributes of a sample, which are key attributes and which are non-key attributes are influenced by the context within the relevant instance. In Bongard-Logo problems, some instances focus on more macroscopic attributes of the image, such as its closure and concavity & convexity; while other instances focus on more detailed attributes, such as the number of entities in the image and the order of their arrangement. This is different from RPM problems, where the key and non-key attributes are consistent across all instances within the RPM.
Given the differences in determining correct solutions, or in identifying key and non-key attributes, between Bongard-Logo and RPM problems, the application of our proposed Funny approach differs when applied to these two types of problems. We will discuss them separately.
VI-A Funny for RPM problems
Given that the key attributes and non-key attributes in the RPM problem are fixed, with key attributes being key across all instances and non-key attributes being unfavored in any instance, this paper proposes that the representations of samples with the same key attributes can be mapped into a single Gaussian distribution. Within this Gaussian distribution, the key attributes of the samples are identical, but the non-key attributes vary. Consequently, the distribution of representations for all samples in the entire RPM problem is formulated as a mixture of Gaussian distributions. By exposing Valen to this mixture of Gaussian distributions instead of individual sample representations, we effectively clarify the distribution of correct solutions for Valen. Our intention is to achieve the transformation of the sample representation space as illustrated in the Figure 15.
Figure 15: Modeling the representation distribution as a mixture of Gaussian distributions.
Therefore, the purpose of the Funny method is to plan the distribution of samples in the RPM problem as a mixture of Gaussian distributions, with the key attributes serving as the component means.
The design of Valen’s Funny method, applied to tackle the RPM problem, is quite straightforward. The Funny method separately extracts the key attributes and non-key attributes of RPM images, treating the key attributes as the center and the non-key attributes as biases relative to this center. Therefore, we refer to the representation of the key attributes extracted from the RPM image $x_{i}$ as $center_{i}$, and the representation of the non-key attributes as $bias_{i}$. Interestingly, the sum of $center_{i}$ and $bias_{i}$ can be viewed as samples drawn from a Gaussian distribution $N(center_{i},\sigma^{2})$, which represents a distribution with a mean of $center_{i}$ and an unknown variance $\sigma^{2}$. When the $bias_{i}$ follow a standard normal distribution, this Gaussian distribution can be considered as $N(center_{i},1)$, which is a distribution with a mean of $center_{i}$ and a variance of 1. In this scenario, representations of samples with the same key attributes but different non-key attributes only differ in their biases. This implies that the Funny method successfully maps representations of samples with the same key attributes into the same Gaussian distribution.
Based on this hypothesis, we design a multi-viewpoint image representation extractor as illustrated in Figure 16, which replaces Valen’s original representation extractor when equipped with the Funny method.
Figure 16: Funny representation extractor
As can be seen in the figure 16, Funny encodes the image sample $x_{i}$ of the reasoning problem into two sets of homomorphic multi-viewpoint representations, named $center_{i}$ and $bias_{i}$ respectively. The final multi-viewpoint representation $r_{i}$ of the image $x_{i}$ extracted by the Funny representation extractor is obtained by summing $center_{i}$ and $bias_{i}$.
However, the new representation extractor of the Funny method, which encodes images $x_{i}$ into $center_{i}$ and $bias_{i}$, still faces two challenges: 1. How to ensure that the $center_{i}$ stores key attributes while the $bias_{i}$ stores non-key attributes? 2. Given that the union of key and non-key attributes constitutes all attributes, how to guarantee that the combination of $center_{i}$ and $bias_{i}$ can store all attributes of the image $x_{i}$? These two questions pose challenges to both the center and bias individually and collectively.
This paper argues that by enforcing $bias_{i}$ to follow a standard normal distribution $N(0,1)$ and simultaneously enforcing all random sampling results, denoted as $\epsilon$, from a Gaussian distribution $N(center_{i},1)$ to assume the same reasoning identity, it can be ensured that $center_{i}$ stores at least the key attributes. This is because, when samples drawn from Gaussian distribution $N(center_{i},1)$ all share the same reasoning identity, $bias_{i}$, which follows a standard normal distribution $N(0,1)$, ceases to influence the reasoning, and $center_{i}$ takes full responsibility for the reasoning task, indicating that $center_{i}$ stores at least the key attributes at this point.
Furthermore, to ensure that the combination of $center_{i}$ and $bias_{i}$ can store all attributes within image $x_{i}$, and consequently guarantee that the $bias_{i}$ stores the image attributes omitted by the $center_{i}$, it suffices to maximize the mutual information between combination of $center_{i}$ and $bias_{i}$ and all attributes within image $x_{i}$. The most direct method to maximize the mutual information between image attributes and image representations is to append an additional reconstruction task. Therefore, when the Valen framework incorporates the Funny method, its structure should be as illustrated in the Figure 17.
Figure 17: The Valen framework incorporates the Funny method
In the figure 17, we can observe that the representation of key attribute, which is $\{center_{i}|i\in[1,9]\}$, extracted by the Funny extractor for a $3\times 3$ image matrix $\{x_{i}|i\in[1,9]\}$ is superimposed with the sampling results $\epsilon$ from a Gaussian normal distribution $N(0,1)$ and then fed into the regular Valen framework to undertake subsequent reasoning tasks. This is done to clarify with the help of Valen that the representations drawn from the Gaussian distribution $N(center,1)$ have the same reasoning identity.
Furthermore, in the figure 17, we can observe that Funny introduces three additional regression tasks based on mean squared error (MSE) loss to Valen, which are indicated by dashed lines in red, yellow, and green. The green dashed line indicates the regression task between the representation of non-key attributes, which is $\{bias_{i}|i\in[1,9]\}$, and the random sampling results $\epsilon$ from the standard normal distribution $N(0,1)$. It is worth noting that the $\epsilon$ here needs to be resampled and cannot be the one previously superimposed on the $\{center_{i}|i\in[1,9]\}$. The objective of this green task is to enforce the bias to conform to the standard normal distribution $N(0,1)$. The red and yellow dashed lines indicate the two regression tasks performed by the Funny decoder. The task indicated by the red dashed line aims to reconstruct the corresponding RPM image $x_{i}$ from the representation obtained by adding $center_{i}$ and $bias_{i}$. The task indicated by the yellow dashed line aims to reconstruct RPM images from the representation obtained by adding $center_{i}$ and $\epsilon$.
The objectives of these two reconstruction tasks are vastly different. The reconstruction task indicated by the red dashed line is aimed at maximizing the conditional distribution $P(x_{i}|center_{i}+bias_{i})$ during the training process of the Funny decoder. On the other hand, the reconstruction task indicated by the yellow dashed line is intended to ensure that the output of the Funny decoder follows the distribution of RPM images, thereby guaranteeing that the Funny decoder can effectively function as a distribution transformation function.
The objective of the red reconstruction task, which is to maximize the conditional distribution $P(x_{i}|center_{i}+bias_{i})$, is equivalent to maximizing the mutual information between the attributes of $x_{i}$ and $center_{i}+bias_{i}$. This mutual information can be calculated using the following formula:
$\text{MI}(x_{i},{center}_{i}+{bias}_{i})=H(x_{i})-H(x_{i}|{center}_{i}+{bias}_{i})$ (2)
where $H(x_{i})$ is the entropy of $x_{i}$, and $H(x_{i}|center_{i}+bias_{i})$ is the conditional entropy of $x_{i}$ given $center_{i}+bias_{i}$. This formula quantifies the amount of information shared between $x_{i}$ and the combination of ${center}_{i}$ and ${bias}_{i}$. Additionally, the operations similar to the yellow reconstruction task, which involve constraining the output of the decoder to follow a certain distribution through MSE-based reconstruction tasks, have already been well-established and widely used in VAE[8].
Overall, Funny introduces three new MSE-based regression tasks for Valen. The green task aims to enforce the bias to conform to the standard normal distribution $N(0,1)$. The red task aims to maximize the mutual information between center+bias and $x_{i}$. The yellow task aims to enforce the output of the Funny decoder to conform to the distribution of $x_{i}$.
VI-B Funny, or not Funny
What is of paramount importance is that, potentially due to the specificity of MSE loss or unknown reasons arising from forcibly combining regression tasks with reasoning tasks, a decoder with a conventional structure cannot bear the aforementioned theoretical framework. If we apply a deconvolution structure to the backbone of the Funny decoder, its optimization becomes extremely challenging, making it difficult to achieve satisfactory regression results and supervise the mutual information between $center_{i}+bias_{i}$ and $x_{i}$. Alternatively, if we replace the Funny decoder with a ViT structure symmetric to the representation extractor, although we obtain excellent regression results, it leads to a decrease in the normal reasoning accuracy of Valen instead of an increase.
Unfortunately and unfunny, there is a disconnect between theory and practice. This paper speculates that this may be due to the impact of the regression task in Funny on the parameter optimization of the reasoning task in Valen. An attempt was made to linearly search for the scaling coefficient between the MSE-based regression task introduced by Funny and the CE-based reasoning task in Valen, but it did not yield an effective solution. Anyway, this paper does not intend to bridge this gap between theory and practice by adjusting the scaling coefficients among the various terms in the loss function.
Fortunately and funny, this paper presents a novel decoder architecture, which we term the “Half-split Decoder” to address this issue. The structure of our designed Half-split Decoder is illustrated in the Figure 18.
Figure 18: Half-split Decoder
The upper half, separated by the blue dashed line in figure 18, represents the representation extractor of Funny, which has been previously shown in Figure 16. The lower half depicts the structure of our designed Half-split Decoder.
In the figure 18, we can observe that our Half-split Decoder is composed of two identical structures, with each structure’s parameters optimized independently without parameter sharing. It is evident that each individual structure resembles an inverted version of a ViT[17], which we will not delve into further in this paper. These two identical structures regress the multi-viewpoint representation $r_{i}$ into two outcomes, $xu_{i}$ and $xd_{i}$, respectively. Here $r_{i}$ can be derived either by summing ${center}_{i}$ with $\epsilon$ or by summing ${center}_{i}$ and $bias_{i}$, corresponding to the two regression requirements of Funny. However, the core of the Half-split Decoder lies not in these structures but in the two distinct activation functions illustrated in the figure 18. We represent these two activation functions as follows:
$f_{1}(xu_{i})=\frac{\text{Sigmoid}(xu_{i})+1}{2},f_{2}(xd_{i})=\frac{\text{Sigmoid}(xd_{i})}{2}$ (3)
In the above equation, it is evident that the value range of $f_{1}(xu_{i})$ lies within $(0.5,1)$, while the value range of $f_{2}(xd_{i})$ belongs to $(0,0.5)$. Although the value ranges of both $f_{1}(xu_{i})$ and $f_{2}(xd_{i})$ differ from the value range of the original RPM image $x_{i}$, which is $(0,1)$, we still compute the MSE loss between $f_{1}(xu_{i})$ and $x_{i}$, as well as between $f_{2}(xd_{i})$ and $x_{i}$. The sum of these two errors is used as the new loss function for the two regression tasks associated with the Funny decoder. This loss function can be represented by the following equation:
$\ell_{\text{Funny}}=\text{MSE}(f_{1}(xu_{i}),x_{i})+\text{MSE}(f_{2}(xd_{i}),x_{i})$ (4)
When both $xu_{i}$ and $xd_{i}$ in the formula are decoded from ${center}_{i}+{bias}_{i}$, the loss function is used to supervise the Mutual Information $\text{MI}(x_{i},{center}_{i}+{bias}_{i})$. Conversely, when both $xu_{i}$ and $xd_{i}$ are decoded from ${center}_{i}+\epsilon$, the loss function is employed to supervise the distribution that the output of the Funny decoder follows.
The design of the new loss function and the new decoder structure is based on a hypothesis in this paper. We hypothesize that the new regression task introduced by Funny will complete ahead of the reasoning task inherent to Valen. This premature completion of the regression task may reject further optimization or adjustment of the multi-viewpoint representations, subsequently making it difficult for these representations to optimize and adapt for the reasoning task. Therefore, we have designed the Half-split Decoder for Funny, with the aim of ensuring that the regression task introduced by Funny remains in an unfinished state from the perspective of the loss function. This ensures that the MSE loss does not decrease to zero in the early stages of optimizing the reasoning task, thereby preventing the multi-viewpoint representations from becoming stagnant.
VI-C Funny for Bongard-Logo problem
Funny is inherently a method that analyzes the key degree of sample attributes and modifies the sample representation space accordingly. Since the key degree of an attribute in Bongard-Logo is determined by contextual information, the previous form of Funny is difficult to directly transfer to the Bongard-Logo problem. Therefore, we innovate the Funny method.
This paper argues that we can seek a series of unlearnable Bongard-Logo image sample mapping functions $F(x_{i})$ that do not alter the key degree of image attributes in the current contextual environment. Additionally, these functions do not change the key attributes of the input image but produce varying degrees of alteration to non-key attributes. Clearly, we have not abandoned the motivation of Funny: to make it explicit to Valen that key attributes determine reasoning identity, while non-key attributes do not affect reasoning identity. The previous form of Funny achieved this by mapping the representations of samples with the same key attributes into a Gaussian distribution with a variance of 1.
So, how do we find such functions? Through Human-in-the-Loop (HITL). HITL refers to situations where human experts or users intervene and provide feedback or guidance during the operation or training of machine learning or deep learning models. By analyzing the three coexisting concepts in Bongard-Logo, we subsequently discovered that some direct data augmentation methods, such as horizontal or vertical flipping, and 90-degree, 180-degree, and 270-degree rotations, can effectively serve as the function $F(x_{i})$. Therefore, when we intend to apply the Funny method to Valen for addressing the Bongard-Logo problem, it suffices to randomly apply the aforementioned data augmentation methods to the training samples $x_{i}$ during the regular training process of Valen. It is noteworthy that, unlike in the Bongard-Logo problem, , these data augmentation methods may disrupt the key attributes in RPM problem. In other words, the Funny method tailored for the Bongard-Logo problem cannot be transferred to RPM problems, and vice versa.
VII SBR: Method for Planning the Distribution of Reasoning Pattens.
Given the success of the Funny method in planning the representation distribution of samples, this paper argues that it is also necessary to plan the representation of the patterns, which is $\{P_{nm}|m\in[1,M]\}_{v}$, extracted by Valen for visual reasoning problems.
However, there are no key or non-key parts in the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ encoded by Valen for RPM problems or Bongard-Logo problems, and even currently, the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ lacks any interpretability. Therefore, our Funny method is difficult to directly apply to the pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$.
Therefore, we have designed a new method, SBR (Supervised Representation Distribution Planning Method). This method can explicitly model the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$ as a mixture of Gaussian distributions. The idea of SBR is straightforward: SBR defines the distribution form of $\{P_{nm}|m\in[1,M]\}_{v}$ as a mixture of Gaussian distributions and provides additional supervisory signals for the mean and variance of each sub-component in this hypothetical distribution.
When establishing a Gaussian Mixture Model (GMM), providing means, covariance matrices, and component weights is a necessary procedure. However, the objective of this paper is to model the distribution of pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$ as a more interpretable and orderly mixture of Gaussian distributions, which involves altering the original form of the $\{P_{nm}|m\in[1,M]\}_{v}$ distribution. This implies that we need to provide a reasonable number of components for the hypothesized distribution form of $\{P_{nm}|m\in[1,M]\}_{v}$, namely the mixture of Gaussian distributions, assign reasonable means and variances to these components, and provide the covariance matrices between the components. Otherwise, forced and incorrect distribution modeling approaches may detrimentally affect the accuracy of Valen’s reasoning.
It is noteworthy that in the RPM problems, each instance is not only designed with problem statements, option pools, and correct option labels, but also accompanied by metadata. This metadata provides a detailed description of the progressive patterns inherent in each instance. Interestingly, the intended meaning of $\{P_{nm}|m\in[1,M]\}_{v}$ encoded by Valen for an RPM instance also corresponds to a progressive pattern.
Therefore, we decide to use metadata to provide relevant information about the target form of the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$. By enumerating all the metadata within the instances of the RPM problem, the enumerated metadata list can offer ideal information about the component centers for the target distribution form of $\{P_{nm}|m\in[1,M]\}_{v}$. These enumeration results are mutually independent, owing to the characteristics of metadata.
In short, we enumerate all the metadata within the instances of the RPM problem to obtain a metadata list. After processing each element in the list as a Gaussian distribution, the entire list forms a mixture of Gaussian distribution. Subsequently, we naturally assign $\{P_{nm}|m\in[1,M]\}_{v}$ to the corresponding component of this mixture model, thereby achieving our goal of planning the distribution of pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$. This process can be visualized as Figure 19.
Figure 19: The framework of SBR
In the figure 19, it can be seen that we assume the number of elements in the metadata list to be $S$. Since there are often multiple decoupled progressive patterns within the RPM matrix, such as the “shape” and “line” progressive patterns present in the PGM problem[16], and the related metadata fully describes these progressive patterns, we decide to dissect all the progressive patterns recorded in the metadata and use them to individually guide the assignment of the $M$ progressive patterns stored in the $\{P_{nm}|m\in[1,M]\}_{v}$ to their respective distribution components. In the figure, we dissect the two progressive patterns, “pattern 3” and “pattern S”, recorded in the metadata and use them separately to guide the distribution attribution of $\{P_{n1}|n\in[1,9]\}_{v}$ and $\{P_{n2}|n\in[1,9]\}_{v}$. Evidently, when Valen is combined with the SBR method, the value of $M$ is no longer an arbitrarily set hyperparameter, but rather needs to be determined based on the types of decoupled progressive patterns stored in the metadata. For example, when the SBR method is applied to the PGM problem, the value of $M$ in Valen needs to be set to 2.
Regarding how we process each element in the enumerated metadata list into distributions, we choose to use a standard Transformer-Encoder to encode the elements, which are descriptions of the progressive patterns, into vectors directly. A list containing $S$ elements will be encoded into $S$ vectors, which we denote as $\{q_{s}|s\in[1,S]\}$ here. If the first half of $q_{s}$ is regarded as the mean of a Gaussian distribution and the second half as the logarithm of variance, then $\{q_{s}|s\in[1,S]\}$ can represent a mixture of Gaussian distributions. We can again utilize the reparameterization technique shown in Figure 14 to reparameterize $\{q_{s}|s\in[1,S]\}$ into a new set of vectors, which we denote as $\{Q_{st}|s\in[1,S],t\in[1,T]\}$. In $\{Q_{st}|s\in[1,S],t\in[1,T]\}$, it can be observed that we reparameterized $\{q_{s}|s\in[1,S]\}$ $T$ times, where $T$ is set to 10 in this paper.
Regarding the process indicated by the light blue and purple lines in the figure 19, where $\{P_{n1}|n\in[1,9]\}_{v}$ and $\{P_{n2}|n\in[1,9]\}_{v}$ are separately compressed into the corresponding Gaussian distribution $q_{3}$ and $q_{S}$, we achieve this by setting an additional loss function. This loss function can be expressed as follows:
$\{\bar{P}_{nm}\}$ $=\frac{1}{V}\sum_{v=1}^{V}\{P_{nm}\}_{v}$ (5)
${\ell_{\text{SBR}}}$ $=-\sum_{m=1}^{M}\sum_{\tilde{s}=1}^{S}\sum_{n=1}^{9}\sum_{t=1}^{T}\text{meta}_{{m\tilde{s}}}\cdot\log\frac{{{e^{({\bar{P}_{nm}}\cdot{Q_{\tilde{s}t}})/\tau}}}}{{\sum\nolimits_{s=1}^{S}{{e^{({\bar{P}_{nm}}\cdot{Q_{{s}t}})/\tau}}}}}$ (6)
Where $\text{meta}_{{m\tilde{s}}}$ represents whether the the $m$-th progressive pattern recorded in the metadata of current instance matches the $\tilde{s}$-th element in the metadata list and $\tau$ is a learnable parameter. As can be seen from the formula (5), considering the overhead of computational resources, we do not individually compress each viewpoint in $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$ into the constructed mixture of Gaussian distributions, but instead first calculate the viewpoint average $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$ of $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$ and then plan the distribution of $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$.
Formula (5) has clearly elucidated the mechanism of SBR. SBR calculates the probability of each of the $M$ pattern representations in $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$ belonging to every component in the mixture of Gaussian distributions $\{q_{s}|s\in[1,S]\}$, thereby obtaining $M\times S$ probability values. Subsequently, SBR applies the CE loss to constrain the $M\times S$ probabilities according to the indications of metadata. These $M\times S$ probabilities are obtained by calculating the sample cosine distance between $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$ and $\{Q_{st}|s\in[1,S],t\in[1,T]\}$, and normalizing these distances through the softmax function.
Unfortunately, the descriptions of clustering patterns in Bongard-Logo are overly simplistic, rendering the SBR a specialized solution tailored specifically for RPM problems.
VIII Experiment
All experiments were programmed in Python using the PyTorch[43] framework.
VIII-A Experiment on RAVEN
Initially, we tested Valen on RAVEN and I-RAVEN datasets to show its potential. For a fair comparison, we used the same settings and equipment as RS-Tran[28] and Triple-CFN[32], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
TABLE I: Reasoning Accuracies on RAVEN and I-RAVEN.
Test Accuracy(%)
Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
SAVIR-T [24] 94.0/98.1 97.8/99.5 94.7/98.1 83.8/93.8 97.8/99.6 98.2/99.1 97.6/99.5 88.0/97.2
SCL [23, 24] 91.6/95.0 98.1/99.0 91.0/96.2 82.5/89.5 96.8/97.9 96.5/97.1 96.0/97.6 80.1/87.7
MRNet [21] 96.6/- -/- -/- -/- -/- -/- -/- -/-
RS-TRAN[28] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
Triple-CFN[32] 98.9/99.1 100.0/100.0 99.7/99.8 96.2/97.5 99.8/99.9 99.8/99.9 99.9/99.9 97.0/97.3
Triple-CFN+Re-space[32] 99.4/99.6 100.0/100.0 99.7/99.8 98.0/99.1 99.9/100.0 99.9/100.0 99.9/99.9 98.5/99.0
Valen 99.5/99.7 100.0/100.0 99.7/99.8 98.2/99.2 99.9/100.0 99.9/100.0 99.9/99.9 98.8/99.2
The experimental records in Table I demonstrate that existing excellent RPM solvers, including Valen, exhibit varying reasoning accuracy on the RAVEN [15] and I-RAVEN [37] datasets. The I-RAVEN dataset was designed to address limitations in the RAVEN’s option pool by enhancing its negative sample setup, such as increasing the number of attributes with shifts and the magnitude of attribute shifts. However, these modifications did not alter the configuration of the primary samples or the distribution of correct solutions. Nonetheless, the same solvers still achieved different reasoning accuracies on RAVEN and I-RAVEN. This observation aligns with our perspective in Tine: the learning objective of probabilistic highlighting solvers is not the distribution of correct solutions, but rather the distribution delineated by both primary and auxiliary samples.
VIII-B Experiment on Bongard-Logo
We conducted experiments on Bongard-Logo. The experimental results are presented in the table II.
TABLE II: Reasoning Accuracies of models on Bongard-Logo.
Accuracy(%)
Model Train FF BA CM NV
MetaOptNet 75.9 60.3 71.6 65.9 67.5
ANIL 69.7 56.6 59.0 59.6 61.0
Meta-Baseline-SC 75.4 66.3 73.3 63.5 63.9
Meta-Baseline-MoCo[42] 81.2 65.9 72.2 63.9 64.7
WReN-Bongard 78.7 50.1 50.9 53.8 54.3
SBSD 83.7 75.2 91.5 71.0 74.1
PMoC 92.0 90.6 97.7 77.3 76.0
Triple-CFN 93.2 92.0 98.2 78.0 78.1
Valen 92.5 91.0 97.8 77.5 76.5
Valen+Tine 93.3 92.4 98.2 78.0 77.8
Funny+Valen 95.1 93.4 98.8 80.3 79.8
Funny+Valen+Tine 96.0 95.2 99.3 84.6 82.9
Observing table II, The performance of Funny, Valen and Tine were promising.
VIII-C Experiment on PGM
In this paper, we conducted experiments related to Valen on PGM. And the results are recorded in Tabel III.
TABLE III: Reasoning Accuracies of models on PGM.
Model Test Accuracy(%)
SAVIR-T [24] 91.2
SCL [23, 24] 88.9
MRNet [21] 94.5
RS-CNN[28] 82.8
RS-TRAN[28] 97.5
Triple-CFN[32] 97.8
Triple-CFN+Re-space layer[32] 98.2
Valen 98.5
Valen+Tine 98.8
Funny+Valen 99.0
RS-Tran+Tranclip[28] 99.0
Meta Triple-CFN[32] 98.4
Meta Triple-CFN+Re-space layer[32] 99.3
SBR+Valen 99.4
From the table III, it can be observed that the SBR, Tine, and Funny methods have all contributed to varying degrees of improvement in Valen’s performance on RPM problems. Valen combined with the Funny method is the first to achieve a reasoning accuracy exceeding 99.0% on the PGM problem without relying on any additional supervisory signals. Furthermore, the integration of SBR with Valen is the first approach to utilize metadata as an additional supervisory signal, achieving a reasoning accuracy of 99.4%. Additionally, SBR provides Valen with interpretability by endowing it with a progressive pattern extraction process. The SBR method presented in this paper is not the first to consider utilizing the accompanying metadata within RPM problem instances to assist the solver in reasoning. Previous works[21] have explored various approaches to leveraging metadata, but they ultimately found that forcibly imposing additional metadata learning tasks on their solvers was counterproductive. RS-Tran[28] and Triple-CFN[28] adopted an indirect manner of utilizing metadata, specifically, RS-Tran pre-trained its representation encoder using metadata, and Triple-CFN required a similar warm-start process. However, the SBR method in this paper is the first to directly utilize metadata to model the distribution of progressive pattern representations, which is an unexplored approach in previous works and has achieved promising results.
We also conducted experiments on the interpolation and extrapolation problems in PGM, and the results are recorded in Table IV. As can be seen from the table IV, the Funny method brings a remarkable improvement to Valen in interpolation.
TABLE IV: Generalization Results of Valen in PGM.
Accuracy(%)
Dataset Valen Valen+Tine Funny+Valen SBR+Valen
Interpolation 81.2 82.8 87.6 92.5
Extrapolation 18.4 18.5 18.5 12.8
IX Conclusion
In conclusion, this paper introduces a novel baseline model, Valen, for solving visual abstract reasoning tasks, particularly Bongard-Logo problems and RPM problems. Valen, a probabilistic-highlighting model, demonstrates remarkable performance in handling both image clustering reasoning and image progression pattern reasoning. The paper further delves into the underlying mechanisms of probability-highlighting solvers, realizing that these solvers approximate solutions to reasoning problem instances as distributions. To bridge the discrepancies between the estimated solution distribution and the true correct solution distribution, the paper introduces three methods: Tine, Funny, and SBR. The Tine method, based on adversarial learning, aims to assist Valen in estimating a solution distribution closer to the correct one. The Funny method models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution. The SBR method further extends this idea by modeling the distribution of progressive patterns representation as a mixture of Gaussian distributions. Overall, the Funny, Tine, and SBR methods significantly improve Valen’s performance, providing new insights and methods for the study of visual abstract reasoning problems.

### v2

@@
-Funny Valen-Tine: Solving visual abstract reasoning problems through defining the solution distribution
+Funny-Valen-Tine: Planning Solution Distribution Enhances Machine Abstract Reasoning Ability
 Abstract
@@
 Furthermore, we designed the SBR method to model the distribution of progressive patterns representation as a mixture of Gaussian distributions.
-Overall, the Funny, Tine, and SBR methods have significantly improved Valen’s performance, providing new ideas and methods for the study of visual abstract reasoning problems.
+Overall, this paper posits that the key to enhancing the ability of solvers to address visual abstract reasoning problems lies in explicitly planning the distribution of solutions to conform more closely to the distribution of correct solution.
 Index Terms:
-Abstract reasoning, Raven’s Progressive Matrices, Bongard-Logo problem, Adversarial Learning, Gaussian Mixture Model.
+Abstract reasoning, Raven’s Progressive Matrices, Bongard-Logo, Adversarial Learning, Gaussian Mixture Model.
 publicationid: pubid: $0$–$00$ 2021 IEEE
@@
 II-A RPM solvers and Bongard-Logo solvers
-There are numerous excellent solvers for RPM (Raven’s Progressive Matrices) and Bongard-Logo problems, among which CoPINet[19], LEN+teacher[18], DCNet[20], NCD[22], SCL[23], SAVIR-T[24], PrAE[25], NVSA[27], ALANS[26], RS-TRAN[28] and CRAB[29] are notable RPM solvers, while SBPDL[30], PMoC[31] and Triple-CFN[32] are outstanding Bongard-Logo solvers.
+There are numerous excellent solvers for RPM and Bongard-Logo problems, among which CoPINet[19], LEN+teacher[18], DCNet[20], NCD[22], SCL[23], SAVIR-T[24], PrAE[25], NVSA[27], ALANS[26], RS-TRAN[28] and CRAB[29] are notable RPM solvers, while SBPDL[30], PMoC[31] and Triple-CFN[32] are outstanding Bongard-Logo solvers.
 Since the method proposed in this paper is relatively independent and does not have a strong dependency on these existing works, the details about them will not be elaborated further.
@@
 Specifically, Bongard-Logo problems can be categorized as image clustering reasoning, while RPM challenges involve image progression pattern reasoning.
-This paper argues that explicitly defining the distribution of solutions for the designed abstract reasoning problem solver is a straightforward and effective approach.
-To this end, we first designed a general solver, Valen, for both RPM and Bongard-Logo problems.
-Additionally, we developed three accompanying methods: Funny, Tine, and SBR.
-These three methods all aim to explicitly define the distribution of correct solutions for Valen through explicit or implicit approaches.
-Here we clarify a definition: in this paper, samples that conform to the underlying patterns within instances of reasoning problems are referred to as “primary samples.” For example, the correct options within RPM problem instances and the samples in the primary group within Bongard-Logo problem instances are considered primary samples.
-Conversely, samples that do not conform to the underlying patterns, such as the incorrect options in RPM problem instances and the samples in the auxiliary group within Bongard-Logo problem instances, are referred to as “auxiliary samples.”
+In this paper, we designed a general solver, Valen, for both RPM and Bongard-Logo problems.
+This paper proposes that to enhance the ability of solvers to address visual abstract reasoning problems, the distribution of solutions provided by the solver should be explicitly planned to better conform to the distribution of correct solutions.
+Therefore, we developed three accompanying methods: Funny, Tine, and SBR.
+These three methods all aim to explicitly plan the distribution of solutions generated by Valen.
+Here we clarify a definition: in this paper, “primary samples” refer to those that align with the core patterns in reasoning problems, like correct RPM options or primary group samples in Bongard-Logo instances.
+Conversely, “auxiliary samples” are those that deviate from these patterns, such as incorrect RPM options or auxiliary group samples in Bongard-Logo problems.
 IV Valen: a new baseline model for RPM and Bongard-Logo problems
-In this section, we introduce Valen (Visual Abstraction Learning Network), our baseline model designed to address both RPM (Raven’s Progressive Matrices), an image progressive pattern reasoning problem, and Bongard-Logo, an image clustering reasoning problem.
+In this section, we introduce Valen (Visual Abstraction Learning Network), our baseline model designed to address both RPM, an image progressive pattern reasoning problem, and Bongard-Logo, an image clustering reasoning problem.
 IV-A Valen targeting RPM problem
 The RPM problem challenges participants to complete a 3×3 matrix of images by choosing answers from a pool of options.
-A correct answer is one that, when placed in the matrix, maintains a consistent and specific image progression pattern that can be deduced from the incomplete matrix, referred to as the problem statement.
-[39].
+A correct answer is one that, when placed in the matrix, maintains a consistent and specific image progression pattern that can be deduced from the incomplete matrix, referred to as the problem statement[36].
 This paper requires a baseline to support our reasoning work.
@@
 Subsequently, these multi-viewpoint representations $\{r_{i}|i\in[1,9]\}_{v}|_{v=1}^{V}$ will be enumerated by Valen to generate all possible incomplete representations matrix.
-The enumeration results are illustrated in the Figure 5, with the last item of the enumeration is the matrix with a incomplete option position.
+The enumeration results are illustrated in the Figure 5.
 Figure 5: Enumeration results of incomplete matrix
-We denote these enumerated multi-viewpoint incomplete representation matrices as $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$, where $n$ represents the index of the incomplete representation matrix, and $i$ still represents the number of the representation within the incomplete matrix.
-Please note the order of the enumeration results set in the figure.
+We denote these enumerated multi-viewpoint incomplete representation matrices as $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$, where $n$ represents the index of the incomplete representation matrix, and $i$ still represents the index of the representation within the incomplete matrix.
+Notably, the last item of the enumeration is the matrix with a incomplete option position, denoted as $\{r_{9i}|n\in[1,9],i\in[1,8]\}_{v}|_{v=1}^{V}$.
 IV-A2 Extraction of progressive patterns from incomplete matrix
@@
 The figure illustrates the process of the Transformer-Encoder handling $\{r_{9i}|i\in[1,8]\}_{v}$, which represents the last item of the enumeration results of incomplete representation matrices from a specific viewpoint $v$.
-The process of handling other incomplete representation matrices $\{r_{ni}|n\in[1,8]i\in[1,9],i\neq n\}_{v}$ is similar to the one shown in the figure, with no essential differences.
+The process of handling other incomplete representation matrices $\{r_{ni}|n\in[1,8],i\in[1,9],i\neq n\}_{v}$ is similar to the one shown in the figure, with no essential differences.
 In the figure 6, we can observe that we use an MLP to extract all row and column progressive information from the incomplete representation matrix $\{r_{9i}|i\in[1,8]\}_{v}$.
@@
 Therefore, we can interpret these solvers as a conditional probability distributions obeyed by abstract reasoning problems, or understand them as providing distributions of solutions to abstract reasoning problems.
-If we use RPM solvers as examples, then these solvers can be abstracted as a conditional probability distribution, denoted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$.
-(Accordingly, when Valen is applied to Bongard-Logo problems, it can be abstracted as a conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{6},\theta)$).
-In this context, the two conditions of the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ are the statement of the reasoning problem $\{x_{i}\}_{i=1}^{8}$ and the optimizable parameters $\theta$ of the solver.
+If we use RPM solvers as examples, then these solvers can be abstracted as a conditional probability distribution, denoted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$.
+(Accordingly, when Valen is applied to Bongard-Logo problems, it can be abstracted as a conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{6};\theta)$).
+In this context, the two conditions of the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ are the statement of the reasoning problem $\{x_{i}\}_{i=1}^{8}$ and the optimizable parameters $\theta$ of the solver.
 Here, the $x_{\alpha}$ refers to the option that is pending verification.
 This conditional distribution implies that, when given a problem statement $\{x_{i}\}_{i=1}^{8}$ and an option $x_{\alpha}$, it is possible to calculate the probability of $x_{\alpha}$ being a solution to $\{x_{i}\}_{i=1}^{8}$.
-The optimization of probability-highlighting solver can be regarded as optimizing the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$.
-A wide range of loss functions can be used for optimization, but they all share the same objective: to assign high probabilities to the primary samples in the probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ and low probabilities to the auxiliary options.
+The optimization of probability-highlighting solver can be regarded as optimizing the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$.
+A wide range of loss functions can be used for optimization, but they all share the same objective: to assign high probabilities to the primary samples in the probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ and low probabilities to the auxiliary options.
 Such an optimization objective makes us realize that the optimization process of the probability-highlighting solver is not aimed at fitting the distribution of correct solutions to visual abstract reasoning problems, but rather is dedicated to fitting a distribution jointly delimited by primary and auxiliary samples.
-To simplify the explanation of our findings, we assume that the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ follows a Gaussian distribution.
-Consequently, we can explicitly identify the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta})$.
-Then, the negative logarithm of the probability density function of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta})$ can be expressed as follows:
-$-\log(P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta}))=\frac{{(x_{a}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(|\,\sigma_{\theta}|\,)$ (1)
+To simplify the explanation of our findings, we assume that the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ follows a Gaussian distribution.
+Consequently, we can explicitly identify the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ as the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$.
+Then, the negative logarithm of the probability density function of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$ can be expressed as follows:
+$-\log(P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta}))=\frac{{(x_{a}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(|\,\sigma_{\theta}|\,)$ (1)
 Evidently, we have omitted the constant term $-1/2\log(2\pi)$.
-In this formula, $\mu_{\theta}$ and $\sigma_{\theta}$ respectively represent the mean and standard deviation of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$.
-Since we have defined the form of distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a Gaussian distribution, we consequently abstract the optimizable parameters $\theta$ into optimizable mean $\mu_{\theta}$ and standard deviation $\sigma_{\theta}$.
-Optimizing the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta})$ using the loss function is equivalent to optimizing the values of $\mu_{\theta}$ and $\sigma_{\theta}$.
+In this formula, $\mu_{\theta}$ and $\sigma_{\theta}$ respectively represent the mean and standard deviation of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$.
+Since we have defined the form of distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ as a Gaussian distribution, we consequently abstract the optimizable parameters $\theta$ into optimizable mean $\mu_{\theta}$ and standard deviation $\sigma_{\theta}$.
+Optimizing the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$ using the loss function is equivalent to optimizing the values of $\mu_{\theta}$ and $\sigma_{\theta}$.
 Our optimization objective is that when primary samples (the correct options in an RPM problem, or the samples within the main group in a Bongard-Logo problem) are plugged into formula (1), we expect to obtain a smaller value.
@@
 Conversely, constraining the return value to achieve a maximum is equivalent to minimizing the value of $\sigma_{\theta}$ in the formula.
-In other words, the probability-highlighting solver positions the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ for the reasoning problem through primary samples, while the standard deviation of the solution distribution is constrained by auxiliary samples.
+In other words, the probability-highlighting solver positions the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ for the reasoning problem through primary samples, while the standard deviation of the solution distribution is constrained by auxiliary samples.
 It appears to be a discouraging reality that the learning objective of a probability-highlighting solver is not to determine the distribution of the correct solutions to reasoning problems, but rather to approximate a distribution that is delimited by both primary and auxiliary samples within those problems.
@@
 VI Funny: Method for Planning the Distribution of Reasoning Problem Samples.
-Based on the analysis in the previous section, auxiliary samples provided in reasoning tasks constrain the standard deviation of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$, while primary samples locate the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$.
+Based on the analysis in the previous section, auxiliary samples provided in reasoning tasks constrain the standard deviation of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$, while primary samples locate the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$.
 The Tine method generates more judicious auxiliary samples to further constrain the standard deviation.
@@
 Given this series of challenges, this paper proposes the Funny (Framework Utilizing Neural Networks for Yielding Image Representation Distribution) method.
-Our Funny method argues that instead of laboriously adopting adversarial learning strategies to make the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ provided by the solver closer to the correct solution distribution, it is more effective to directly specify the range of correct solutions or, alternatively, to explicitly define the details of the correct solution distribution.
-This paper will next analyze the setting and determination of correct solutions for image progressive pattern reasoning problems such as RPM and image clustering reasoning problems such as Bongard-Logo.
+Our Funny method argues that instead of laboriously adopting adversarial learning strategies to make the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ provided by the solver closer to the correct solution distribution, it is more effective to directly specify the range of correct solutions or, alternatively, to explicitly define the details of the correct solution distribution.
+This paper will next analyze the definition of correct solutions for image progressive pattern reasoning problems such as RPM and image clustering reasoning problems such as Bongard-Logo.
 According to the design of RPM problems, the attributes of an image within the RPM problem are divided into key attributes related to reasoning and non-key attributes unrelated to reasoning.
@@
 These two questions pose challenges to both the center and bias individually and collectively.
-This paper argues that by enforcing $bias_{i}$ to follow a standard normal distribution $N(0,1)$ and simultaneously enforcing all random sampling results, denoted as $\epsilon$, from a Gaussian distribution $N(center_{i},1)$ to assume the same reasoning identity, it can be ensured that $center_{i}$ stores at least the key attributes.
+This paper argues that by enforcing $bias_{i}$ to follow a standard normal distribution $N(0,1)$ and simultaneously enforcing all random sampling results from a Gaussian distribution $N(center_{i},1)$ to assume the same reasoning identity, it can be ensured that $center_{i}$ stores at least the key attributes.
 This is because, when samples drawn from Gaussian distribution $N(center_{i},1)$ all share the same reasoning identity, $bias_{i}$, which follows a standard normal distribution $N(0,1)$, ceases to influence the reasoning, and $center_{i}$ takes full responsibility for the reasoning task, indicating that $center_{i}$ stores at least the key attributes at this point.
@@
 Additionally, the operations similar to the yellow reconstruction task, which involve constraining the output of the decoder to follow a certain distribution through MSE-based reconstruction tasks, have already been well-established and widely used in VAE[8].
-Overall, Funny introduces three new MSE-based regression tasks for Valen.
-The green task aims to enforce the bias to conform to the standard normal distribution $N(0,1)$.
-The red task aims to maximize the mutual information between center+bias and $x_{i}$.
-The yellow task aims to enforce the output of the Funny decoder to conform to the distribution of $x_{i}$.
-VI-B Funny, or not Funny
-What is of paramount importance is that, potentially due to the specificity of MSE loss or unknown reasons arising from forcibly combining regression tasks with reasoning tasks, a decoder with a conventional structure cannot bear the aforementioned theoretical framework.
+Our design intention for Funny is quite clear.
+We hope that, with the assistance of Valen, Funny can compel $center_{i}$ to store more key attributes related to reasoning, while $bias_{i}$ accumulates more non-key attributes unrelated to reasoning that could potentially interfere with it.
+By doing so, we aim to model the distribution of multi-viewpoint representations, generated by the $center_{i}+bias_{i}$, as a mixed Gaussian distribution that aids Valen in its reasoning process.
+The relationship between Funny and Valen is mutually beneficial and symbiotic.
+Although subsequent experiments demonstrated that Funny is a significantly effective method, given the fact that obtaining disentangled representations under unsupervised learning is not achievable [40], this paper holds a pessimistic attitude towards the extent of Funny’s decoupling of key and non-key attributes.
+VI-B Funny not funny
+What is of paramount importance is that, a Funny decoder with a conventional structure cannot bear the aforementioned theoretical framework.
 If we apply a deconvolution structure to the backbone of the Funny decoder, its optimization becomes extremely challenging, making it difficult to achieve satisfactory regression results and supervise the mutual information between $center_{i}+bias_{i}$ and $x_{i}$.
 Alternatively, if we replace the Funny decoder with a ViT structure symmetric to the representation extractor, although we obtain excellent regression results, it leads to a decrease in the normal reasoning accuracy of Valen instead of an increase.
-Unfortunately and unfunny, there is a disconnect between theory and practice.
+Unfortunately and unfunny, there is a gep between theory and practice.
 This paper speculates that this may be due to the impact of the regression task in Funny on the parameter optimization of the reasoning task in Valen.
+If we choose the normal ViT as the Funny decoder, the introduced regression task will be completed before the Valen’s reasoning task, because the optimization speed of the regression task is faster than that of the reasoning task at this time.
+We posit that the premature completion of the regression task may result in the $center_{i}$ storing more non-key attribute.
+Because we have set up the regression task from $center_{i}+\epsilon$ to $x_{i}$.
+This paper speculates that this might be the issue: when the regression task approaches completion prematurely, it may lead to $center_{i}$ falling into the gradient vanishing zone, subsequently rendering it difficult for $center_{i}$ to make further optimizations for the reasoning task.
+The non-key attribute stored in $center_{i}$ can disrupt the reasoning process of Valen.
 An attempt was made to linearly search for the scaling coefficient between the MSE-based regression task introduced by Funny and the CE-based reasoning task in Valen, but it did not yield an effective solution.
@@
 Conversely, when both $xu_{i}$ and $xd_{i}$ are decoded from ${center}_{i}+\epsilon$, the loss function is employed to supervise the distribution that the output of the Funny decoder follows.
-The design of the new loss function and the new decoder structure is based on a hypothesis in this paper.
-We hypothesize that the new regression task introduced by Funny will complete ahead of the reasoning task inherent to Valen.
-This premature completion of the regression task may reject further optimization or adjustment of the multi-viewpoint representations, subsequently making it difficult for these representations to optimize and adapt for the reasoning task.
-Therefore, we have designed the Half-split Decoder for Funny, with the aim of ensuring that the regression task introduced by Funny remains in an unfinished state from the perspective of the loss function.
-This ensures that the MSE loss does not decrease to zero in the early stages of optimizing the reasoning task, thereby preventing the multi-viewpoint representations from becoming stagnant.
+This Half-split Decoder can ensure that the loss of the regression task introduced by Funny does not decrease to zero in the early stage of the reasoning task of Valen, thereby preventing the multi-view representation from becoming stagnant.
 VI-C Funny for Bongard-Logo problem
@@
 Therefore, when we intend to apply the Funny method to Valen for addressing the Bongard-Logo problem, it suffices to randomly apply the aforementioned data augmentation methods to the training samples $x_{i}$ during the regular training process of Valen.
-It is noteworthy that, unlike in the Bongard-Logo problem, , these data augmentation methods may disrupt the key attributes in RPM problem.
+It is noteworthy that, unlike in the Bongard-Logo problem, these data augmentation methods may disrupt the key attributes in RPM problem.
 In other words, the Funny method tailored for the Bongard-Logo problem cannot be transferred to RPM problems, and vice versa.
 VII SBR: Method for Planning the Distribution of Reasoning Pattens.
-Given the success of the Funny method in planning the representation distribution of samples, this paper argues that it is also necessary to plan the representation of the patterns, which is $\{P_{nm}|m\in[1,M]\}_{v}$, extracted by Valen for visual reasoning problems.
+This paper suggests that we should also plan the distribution of the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ extracted by Valen for visual abstract reasoning problems, so that it is closer to the distribution of the correct solution’s pattern.
 However, there are no key or non-key parts in the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ encoded by Valen for RPM problems or Bongard-Logo problems, and even currently, the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ lacks any interpretability.
@@
 Funny+Valen 95.1 93.4 98.8 80.3 79.8
-Funny+Valen+Tine 96.0 95.2 99.3 84.6 82.9
+Funny+Valen+Tine 96.0 95.2 99.6 84.6 82.9
 Observing table II, The performance of Funny, Valen and Tine were promising.
@@
 Previous works[21] have explored various approaches to leveraging metadata, but they ultimately found that forcibly imposing additional metadata learning tasks on their solvers was counterproductive.
-RS-Tran[28] and Triple-CFN[28] adopted an indirect manner of utilizing metadata, specifically, RS-Tran pre-trained its representation encoder using metadata, and Triple-CFN required a similar warm-start process.
+RS-Tran[28] and Triple-CFN[28] adopted an indirect manner of utilizing metadata.
+Specifically, RS-Tran utilizes Tranclip[28] to pre-train its representation encoder using metadata, and Triple-CFN required a similar warm-start process.
 However, the SBR method in this paper is the first to directly utilize metadata to model the distribution of progressive pattern representations, which is an unexplored approach in previous works and has achieved promising results.
@@
 The SBR method further extends this idea by modeling the distribution of progressive patterns representation as a mixture of Gaussian distributions.
-Overall, the Funny, Tine, and SBR methods significantly improve Valen’s performance, providing new insights and methods for the study of visual abstract reasoning problems.
+This paper demonstrates the importance of explicitly planning the distribution of solutions in solving visual abstract reasoning problems through the successful application of the SBR, Funny, Valen and Tine.

### v3

@@
 Abstract
-The importance of visual abstruct reasoning problems in the field of image processing is beyond words.
+The importance of visual abstract reasoning problems in the field of image processing cannot be overstated.
 Both Bongard-Logo problems and Raven’s Progressive Matrices (RPM) belong to the domain of visual abstract reasoning tasks, with Bongard-Logo categorized as image clustering reasoning and RPM involving image progression pattern reasoning.
-This paper introduces a novel baseline model, Valen, which falls under the umbrella of probabilistic highlighting models.
+This paper introduces a novel baseline model, Valen, which falls under the umbrella of probability-highlighting models.
 Valen demonstrates remarkable performance in solving both RPM and Bongard-Logo problems, offering a versatile solution for these reasoning tasks.
-Our investigation extends beyond the application of Valen, delving into the underlying mechanisms of probability-highlighting solvers when tackling visual abstract reasoning problems.
-We revisit how these solvers handle instances in tasks like RPM and Bongard-Logo, realizing that they approximate the solution to each reasoning problem instance as a distribution delineated by primary and auxiliary samples.
-This led us to propose that the learning objective of probability-highlighting solvers is not the distribution of correct solutions but rather a distribution defined by both primary and auxiliary samples.
-To bridge the discrepancies, we introduced the Tine method, an adversarial learning-based approach designed to assist Valen in estimating a solution distribution closer to the correct one.
-However, Tine’s adversarial framework also introduces issues such as unstable training.
-Reflecting on Tine, we propose modeling the sample distribution of reasoning problems as a mixture of Gaussian distributions, leading to the development of the Funny method.
-Funny effectively models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution.
-Furthermore, we designed the SBR method to model the distribution of progressive patterns representation as a mixture of Gaussian distributions.
-Overall, this paper posits that the key to enhancing the ability of solvers to address visual abstract reasoning problems lies in explicitly planning the distribution of solutions to conform more closely to the distribution of correct solution.
-Index Terms:
-Abstract reasoning, Raven’s Progressive Matrices, Bongard-Logo, Adversarial Learning, Gaussian Mixture Model.
-publicationid: pubid: $0$–$00$ 2021 IEEE
+Our investigation extends beyond the application of Valen, delving into the underlying mechanisms of probability-highlighting solvers.
+In revisiting how these solvers handle RPM and Bongard-Logo tasks, we realize that they approximate the solution to each reasoning problem as a distribution in which primary samples are compliant while auxiliary samples are not.
+This prompts us to propose that the learning objective of probability-highlighting solvers is not the distribution of correct solutions but rather one jointly delineated by primary and auxiliary samples.
+To bridge the discrepancies, we introduced the Tine method, an adversarial learning-based approach that helps Valen estimate a distribution close to that of the correct solutions.
+However, adversarial training in Tine suffers from instability.
+Motivated by this limitation, we model the sample distribution of reasoning problems as a mixture of Gaussian distributions, enabling Valen to capture the correct solution distribution more efficiently.
+This non-adversarial methodology leads to the development of the Funny method.
+Building on a similar Gaussian-mixture paradigm, we further propose the SBR method to plan the distribution of progressive pattern representations.
+Overall, this paper contends that the key to enhancing solvers’ ability to address visual abstract reasoning problems lies in explicitly planning the distribution of predicted solutions to approach the correct solution distribution.
+Codes are available in: https://github.com/Yuanbeiming/Funny-Valen-Tine-Planning-Solution-Distribution-Enhances-Machine-Abstract-Reasoning-Ability
 I Introduction
-Deep neural networks have made remarkable achievements in various fields, including computer vision[1, 2, 3], natural language processing[4, 5, 6], generative model[7, 8, 9], visual question answering[10, 11], and visual abstract reasoning[14, 12, 13].
-Visual abstract reasoning plays a pivotal role in the field of image processing.
-It is not only one of the advanced tasks in image processing, requiring systems to recognize objects in images and understand the relationships between them for logical inference, but also a key driver for technological innovation and application expansion in image processing.
-By addressing visual abstract reasoning problems, new algorithms and techniques such as deep learning and graph convolutional networks can be explored and applied, thereby propelling extensive applications and developments in fields like autonomous driving, medical diagnosis, and security surveillance.
-However, this field still faces challenges such as insufficient and imbalanced data, as well as limited model interpretability.
-Future research will focus more on enhancing the robustness, real-time performance, and interpretability of algorithms to further elevate the potential and value of visual abstract reasoning technology.
+Deep neural networks have made remarkable achievements in various fields, including computer vision [1, 2, 3, 4], natural language processing [6, 7, 5, 8], generative models [9, 10, 11], visual question answering tasks [12], and visual abstract reasoning [15, 13, 14].
+Visual abstract reasoning plays a pivotal role in image processing.
+This task not only requires systems to detect objects and understand their relationships for logical inference, but also drives technological innovation and application expansion.
+By tackling such reasoning problems, we can stimulate the development of algorithms and technologies such as deep learning and graph convolutional networks, thereby advancing autonomous driving, medical diagnosis, security surveillance, and related fields.
+However, the field still faces insufficient and imbalanced data, limited model performance, and poor interpretability.
+Even large language models (LLMs), which have delivered breakthroughs in many domains, exhibit a significant performance drop on visual abstract reasoning tasks.
+Future research should therefore prioritize enhancing the abstract reasoning capabilities of machine intelligence and actively use such problems to evaluate the level of machine intelligence.
 The RPM problem [15] and the Bongard problem [13, 14] are two renowned examples within the realm of visual abstract reasoning. These two problems are the target problems that this paper aims to solve.
 I-A RAVEN and PGM
-RPM[14], or Raven’s Progressive Matrices, is a classic visual reasoning intelligence test typically presented as a 3x3 matrix.
-It requires test-takers to observe and analyze visual features such as shape, size, direction, position, and their relationships within the problem statement, identify patterns or rules, and select an answer from given option pool that completes the matrix.
-Due to its non-verbal nature and value in assessing general intelligence, RPM is widely used in psychology, education, neuroscience, and other fields, serving as a tool for evaluating intelligence levels and studying human cognitive processes, intellectual development, and brain function.
-With the advancement of AI technology, researchers are exploring computer algorithms to solve RPM problems, albeit facing challenges.
-RAVEN[15] and PGM[16] are classic examples of RPM problems, with Figure 1 illustrating these two types of problems respectively.
+RPM[15], or Raven’s Progressive Matrices, is a classic visual reasoning intelligence test typically presented as a $\times3$ matrix.
+Test-takers must observe and analyze visual features such as shape, size, direction, position, and their relations, identify the underlying rules, and select the missing element from the given option pool that completes the matrix.
+Thanks to its non-verbal format and proven utility in assessing general intelligence, RPM is widely used in psychology, education, and neuroscience as a tool for evaluating intellectual ability and investigating human cognitive development, brain function, and related processes.
+With advances in AI, researchers are now exploring algorithmic solutions to the RPM problem, though significant challenges remain RAVEN[16] and PGM[17] are classic RPM benchmarks; Figure 1 illustrates sample instances.
+Intuitively, given the remarkable success of LLMs in many domains, one might expect them to handle such problems effortlessly.
+However, a series of studies [20, 18, 19] has shown that LLMs fall far short on RPM tasks, underscoring the unique value of these benchmarks and indicating that abstract reasoning remains a scarce and essential capability for machine intelligence.
 Figure 1: RAVEN instance (a) and PGM instance (b)
 I-B Bongrad-Logo
-Distinct from the image progressive pattern reasoning problems like RPM, the Bongard problem [12, 13] belongs to the category of image clustering reasoning problems.
-It typically involves multiple images divided into primary and auxiliary groups, with images in the primary group sharing an abstract concept governed by a specific rule and those in the auxiliary group deviating from these rules.
-Solving Bongard problems requires deep learning algorithms to accurately classify ungrouped images.
+Distinct from RPM tasks, the Bongard problem [13, 14] belongs to image clustering [38] reasoning problems.
+It typically involves multiple images divided into primary and auxiliary groups: images in the primary group share an abstract concept governed by specific rules, whereas images in the auxiliary group violate these rules.
+Solving Bongard problems demands that deep learning algorithms accurately classify the remaining ungrouped images.
 As a specific instance within abstract reasoning, the Bongard-Logo problem [14] poses a significant challenge due to its inferential complexity.
-Each case comprises 14 images: six from the primary group, six from the auxiliary group, and two for classification.
+Each case comprises 14 images: six from the primary group, six from the auxiliary group, and two ungrouped images to be classified.
 The grouping criterion relies on geometric shapes and their arrangement.
-Figure 2 illustrates a Bongard-Logo case.
+Figure 2 illustrates one such case.
 Figure 2: Bongard-Logo instance
+Bongard-logo problems are categorized into four sets: The Free-form Shape Test Set (FF) aims to evaluate a model’s ability to generalize to more complex free-form shape concepts.
+The Basic Shape Test Set (BA) assesses the model’s generalization capability to new combinations of basic shape concepts.
+The Combinatorial Abstract Shape Test Set (CM) tests the model’s abstract understanding of novel attribute combinations, where all individual attributes have been seen in the training set, but the combinations are entirely new.
+The Novel Abstract Shape Test Set (NV) evaluates the model’s ability to discover new abstract concepts, particularly by withholding one attribute and all its combinations to examine the model’s extrapolation capability.
+Solvers are trained on the training set and tested on these four sets to evaluate their intelligence levels from multiple dimensions.
 II Related work
-II-A RPM solvers and Bongard-Logo solvers
-There are numerous excellent solvers for RPM and Bongard-Logo problems, among which CoPINet[19], LEN+teacher[18], DCNet[20], NCD[22], SCL[23], SAVIR-T[24], PrAE[25], NVSA[27], ALANS[26], RS-TRAN[28] and CRAB[29] are notable RPM solvers, while SBPDL[30], PMoC[31] and Triple-CFN[32] are outstanding Bongard-Logo solvers.
-Since the method proposed in this paper is relatively independent and does not have a strong dependency on these existing works, the details about them will not be elaborated further.
+II-A RPM solver
+Discriminative models for image reasoning, such as CoPINet [22], LEN+teacher [23], and DCNet [24], offer diverse solutions, focusing on learning differences and latent rules.
+Connectionist models: SCL [26], SAVIR-T [27], and neural-symbolic systems: PrAE [28], NVSA [29], ALANS [30] incorporate various techniques to improve reasoning accuracy and interpretability.
+By combining effective methods, RS-TRAN [32] have achieved impressive results on the RPM problem.
+Triple-CFN implicitly extracts information about concepts and reasoning units, and “indexes” this reasoning-unit information according to the concepts, achieving noteworthy reasoning accuracy.
+CRAB [33], based on a Bayesian modelling approach, has established a highly customised, controlled sandbox that aligns seamlessly with its proprietary methodology, embodied in the creation of a new RAVEN database.
+This bespoke environment, although it forfeits inherent core challenges of RAVEN such as solution diversity and uncertainty, has enabled CRAB to achieve remarkable breakthroughs.
+The community looks forward to assessing the broader implications of this approach for future research.
+DRNet [34] employs parallel CNN and ViT to extract local and spatial features, and uses a learnable fusion module to model abstract rules explicitly, offering a concise new framework for visual abstract reasoning.
+NCD [31] converts the unsupervised RPM task into contrastive learning via pseudo labels, and employs negative answer sampling plus decentralized features to adaptively cancel noise and inter-problem centroid shifts, enabling automatic rule representation extraction without any annotation and offering a scalable new paradigm for unsupervised abstract reasoning.
+II-B Bongard solver
+Researchers have tackled Bongard problems along three main avenues: symbolic models, convolutional neural networks, and synthetic-data strategies.
+Depweg et al.
+combined a formal language with symbolic vocabulary and Bayesian inference, yet their method struggled with complex instances and scaled poorly [13].
+Kharagorgiev and Yun pre-trained CNNs on image datasets for feature extraction [35].
+Nie et al.
+experimented with CNN-based meta-learning, but results remained sub-optimal [14].
+These approaches exhibit distinct strengths and weaknesses, underscoring the need for further research.
+Yuan proposed PMoC [36], a probabilistic model that measures how likely auxiliary group samples fit the main group’s distribution—an approach that has shown strong performance.
+Notably, Triple-CFN [37] has successfully applied a unified framework to both Bongard-Logo and RPM problems.
 III Methodology
@@
 Specifically, Bongard-Logo problems can be categorized as image clustering reasoning, while RPM challenges involve image progression pattern reasoning.
-In this paper, we designed a general solver, Valen, for both RPM and Bongard-Logo problems.
-This paper proposes that to enhance the ability of solvers to address visual abstract reasoning problems, the distribution of solutions provided by the solver should be explicitly planned to better conform to the distribution of correct solutions.
-Therefore, we developed three accompanying methods: Funny, Tine, and SBR.
-These three methods all aim to explicitly plan the distribution of solutions generated by Valen.
-Here we clarify a definition: in this paper, “primary samples” refer to those that align with the core patterns in reasoning problems, like correct RPM options or primary group samples in Bongard-Logo instances.
-Conversely, “auxiliary samples” are those that deviate from these patterns, such as incorrect RPM options or auxiliary group samples in Bongard-Logo problems.
+We designed a general solver, Valen, for both RPM and Bongard-Logo problems.
+Furthermore, to enhance a solver’s ability to tackle visual abstract reasoning tasks, we propose that its solution distribution should be explicitly planned to conform more closely to that of correct solutions.
+Therefore, we developed three methods to perform this explicit planning: Funny, Tine and SBR.
+Here we clarify the definitions: “primary samples” are those conforming to the core patterns of reasoning problems (e.g., correct RPM options or primary group samples in Bongard-Logo instances); “auxiliary samples” are those that do not (e.g., incorrect RPM options or auxiliary group samples in Bongard-Logo problems).
 IV Valen: a new baseline model for RPM and Bongard-Logo problems
-In this section, we introduce Valen (Visual Abstraction Learning Network), our baseline model designed to address both RPM, an image progressive pattern reasoning problem, and Bongard-Logo, an image clustering reasoning problem.
+In this section, we introduce Valen (Visual Abstraction Learning Network), our baseline model designed to address both RPM, an image progression pattern reasoning problem, and Bongard-Logo, an image clustering reasoning problem.
 IV-A Valen targeting RPM problem
-The RPM problem challenges participants to complete a 3×3 matrix of images by choosing answers from a pool of options.
-A correct answer is one that, when placed in the matrix, maintains a consistent and specific image progression pattern that can be deduced from the incomplete matrix, referred to as the problem statement[36].
-This paper requires a baseline to support our reasoning work.
-This baseline is expected to determine the plausibility of an RPM matrix based on the progression patterns exhibited within it and assign a probability to reflect this plausibility.
-In RPM problems, there exists a unique design where an incomplete RPM matrix (missing one image) can still convey a complete progression pattern.
-This design is essential as RPM serves as an assessment tool.
-Hence, this paper asserts that the reasoning process of a successful RPM problem solver should be schematically represented as illustrated in the Figure 3.
-Figure 3: The reasoning process of a envisioned RPM problem solver
+The RPM problem challenges participants to complete a $\times3$ image matrix by selecting an answer from an option pool.
+The correct option is the one that preserves the progression patterns reasoned from the incomplete matrix (i.e., the problem statement) [41].
+To this end, we introduce a baseline named Valen: it assesses the plausibility of the completed $\times3$ matrix based on its internal progressive patterns and outputs a corresponding probability.
+A defining feature of RPM problems is that an incomplete matrix (missing one image) still retains sufficient information to reason about the complete progressive patterns.
+Based on this property, we argue that the reasoning process of an effective RPM problem solver should align with the illustration in Figure 3.
+Figure 3: The reasoning process of an effective RPM problem solver
 As seen in the figure, our envisioned RPM solver consists of two modules: the Progressive Pattern Extractor and the Pattern Consistency Analyzer.
 Additionally, our envisioned RPM solver can perform the following three reasoning processes: 1.
-The process indicated by the blue line: When the RPM problem statement is completed by an option from the option pool to form a full $3\times 3$ progression matrix, the solver can enumerate all $N$ types of incomplete matrix of this $3\times 3$ matrix.
+The process indicated by the blue line: When an option from the pool completes the RPM problem statement into a full $3\times 3$ progression matrix, the solver can enumerate all $N$ possible incomplete versions of the completed $\times3$ matrix.
 2.
-The process denoted by the green line: For each enumerated incomplete matrix, the Progressive Pattern Extractor extracts the progression pattern, resulting in a total of $N$ sets of progression patterns.
+The process denoted by the green line: For each enumerated incomplete matrix, the Progressive Pattern Extractor extracts its progressive pattern, yielding $N$ sets of progressive patterns.
 3.
-The process marked by the orange line: The solver utilizes its Pattern Consistency Analyzer to evaluate the consistency of these $N$ sets of progression patterns and outputs a score to quantify this consistency.
+The process marked by the orange line: The solver utilizes its Pattern Consistency Analyzer to evaluate the consistency of these $N$ sets of progressive patterns and outputs a score to quantify this consistency.
 This score also serves as the score for the corresponding option.
-The aforementioned three processes will also constitute the procedure of Valen.
-Since this paper provides a clear definition of the reasoning process for a successful RPM solver, our designed Valen is derived by refining these three processes.
-Specifically, our Valen also consists of three processes: representation extraction of matrix and enumeration of incomplete matrix, extraction of progressive patterns from incomplete matrix, and consistency analysis of progressive patterns, which correspond to the aforementioned three processes respectively.
-We will introduce these three refined processes one by one.
-IV-A1 Representation extraction of matrix and enumeration of incomplete matrix
-In this paper, this process in Valen is designed as the structure shown in Figure 4.
+Since this paper provides a clear definition of the reasoning process for a successful RPM solver, the Valen model is derived by refining these three steps.
+Specifically, Valen comprises three stages: (i) representation extraction of matrices and enumeration of incomplete matrices, (ii) extraction of progressive patterns from incomplete matrices, and (iii) consistency analysis of progressive patterns.
+We describe each stage individually.
+IV-A1 Representation extraction of matrices and enumeration of incomplete matrices
+This process in Valen follows the structure illustrated in Figure 4.
 Figure 4: Representation extraction of matrix
-As shown in figure 4, when the problem statement of the RPM task is completed by a specific option to form a $3\times 3$ progression matrix, denoted as $\{x_{i}|i\in[1,9]\}$, Valen utilizes a basic Vision Transformer (ViT) as a feature extractor to obtain multi-viewpoint representations of $\{x_{i}|i\in[1,9]\}$.
-These representations are denoted as $\{r_{i}|i\in[1,9]\}_{v}$, where $v\in[1,V]$ and $V$ represents the number of viewpoints, $v$ represents the index of the viewpoint, and $i$ represents the index of the image in the matrix.
-The subsequent processes in Valen will handle these multi-viewpoint representation matrices in parallel and equally.
+As shown in this figure, when the problem statement of the RPM task is completed by a specific option to form a $3\times 3$ progression matrix, denoted as $\{x_{i}|i\in[1,9]\}$, Valen utilizes a basic Vision Transformer (ViT) as a feature extractor to obtain multi-viewpoint representations of $\{x_{i}|i\in[1,9]\}$.
+These representations are denoted as $\{r_{i}|i\in[1,9]\}_{v}$, where $v\in[1,V]$, $V$ is the total number of viewpoints, $v$ represents the index of the viewpoint, and $i$ represents the index of the image in the matrix.
+The subsequent processes in Valen handle these multi-viewpoint representation matrices in parallel and on an equal footing.
 The method of processing RPM images into multi-viewpoint representations has been applied in previous works [32, 27].
-Subsequently, these multi-viewpoint representations $\{r_{i}|i\in[1,9]\}_{v}|_{v=1}^{V}$ will be enumerated by Valen to generate all possible incomplete representations matrix.
-The enumeration results are illustrated in the Figure 5.
+Subsequently, these multi-viewpoint representations $\{r_{i}|i\in[1,9]\}_{v}|_{v=1}^{V}$ will be enumerated by Valen to generate all possible incomplete representation matrices.
+The enumeration results are illustrated in Figure 5.
 Figure 5: Enumeration results of incomplete matrix
 We denote these enumerated multi-viewpoint incomplete representation matrices as $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$, where $n$ represents the index of the incomplete representation matrix, and $i$ still represents the index of the representation within the incomplete matrix.
-Notably, the last item of the enumeration is the matrix with a incomplete option position, denoted as $\{r_{9i}|n\in[1,9],i\in[1,8]\}_{v}|_{v=1}^{V}$.
-IV-A2 Extraction of progressive patterns from incomplete matrix
+Notably, the last item of the enumeration is the matrix with an incomplete option position, denoted as $\{r_{9i}|i\in[1,8]\}_{v}|_{v=1}^{V}$.
+IV-A2 Extraction of progressive patterns from incomplete matrices
 In Valen, this process is undertaken by a Transformer-Encoder.
@@
 This asymmetric seq-to-seq implementation (four-to-$M$) has been introduced in related techniques of ViT[21].
-After processing all the incomplete representation matrices $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$ using the standard Transformer-Encoder, we obtain all multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$.
+After processing all the incomplete representation matrices $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$ using the Transformer-Encoder, we obtain all multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$.
 IV-A3 Consistency analysis of progressive patterns
-In this process, Valen will evaluate the consistency among these multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$.
-The task of this consistency analysis will be accomplished by the Transformer-Decoder.
-The process of evaluating the consistency of the progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$ from a specific viewpoint $v$ can be conducted using a Transformer-Decoder, as illustrated in Figure 7.
+In this process, Valen evaluates the consistency among these multi-viewpoint progressive pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$.
+As illustrated in Figure 7, Valen conducts this consistency evaluation among the representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$ from each specific viewpoint $v$ using a Transformer-Decoder.
 Figure 7: Evaluating the consistency of the progressive patterns representation $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$
-In the figure 7, we can observe that the progressive pattern representations $\{P_{9m}|m\in[1,M]\}_{v}$ extracted from the incomplete representation matrix $\{r_{9i}|i\in[1,8]\}_{v}$ are used as queries, while the other progressive pattern representations $\{P_{nm}|n\in[1,8],m\in[1,M]\}_{v}$, which are extracted from the incomplete representation matrices $\{r_{ni}|n\in[1,8],i\in[1,9],i\neq n\}_{v}$, serve as key-value pairs input into the Transformer-Decoder to compute the output vector set.
-The $M$ vectors within the output vector set are then individually mapped to a score by a new MLP.
-These mapped $M$ scores are summed to obtain the consistency score of the progressive patterns under viewpoint $v$.
-Subsequently, We calculate the average of the progressive pattern consistency scores obtained from all viewpoints to determine the overall viewpoint score for the corresponding option.
-Finally, we employ cross-entropy as the loss function to constrain the scores of all options according to the logic of the RPM task, in order to optimize the learnable parameters involved in the aforementioned three processes.
+In the Figure 7, the progressive pattern representations $\{P_{9m}|m\in[1,M]\}_{v}$ extracted from the incomplete representation matrix $\{r_{9i}|i\in[1,8]\}_{v}$ are used as queries, while the remaining progressive pattern representations $\{P_{nm}|n\in[1,8],m\in[1,M]\}_{v}$, which are extracted from the incomplete representation matrices $\{r_{ni}|n\in[1,8],i\in[1,9],i\neq n\}_{v}$, act as key-value pairs for the Transformer-Decoder to compute the output vector set.
+The $M$ output vectors are then individually mapped to scores by an MLP and summed to yield the consistency score of the progressive patterns under viewpoint $v$.
+Subsequently, we compute the average of the progressive pattern consistency scores across all viewpoints to obtain the overall viewpoint score for the corresponding option.
+Finally, cross-entropy loss is employed to align the score distribution with the ground-truth option, thereby optimizing all learnable parameters involved in the aforementioned three processes.
 IV-B Valen targeting Bongard-Logo problem
-We consider Valen as a good baseline model for dealing with image progressive pattern reasoning problems, but it can also handle image clustering reasoning problems such as Bongard-Logo.
-To adapt Valen to the Bongard-Logo problem, we need to transform this problem accordingly.
-In a particular instance of the Bongard-Logo problem, we use $x_{i}$ to represent all the images.
-Specifically, the images in the primary group of this instance are denoted by $\{x_{i}|i\in[1,6]\}$, while those in the auxiliary group are represented by $\{x_{i}|i\in[8,13]\}$.
-Furthermore, $x_{7}$ represents the test sample that should be assigned to the primary group, and $x_{14}$ represents the test sample that should be assigned to the auxiliary group.
-This paper proposes that we can adopt different perspectives to view image clustering reasoning problems, such as the Bongard-Logo problem.
-Specifically, we can treat them as an option evaluation problem akin to the RPM problem.
-To elaborate, the images in the primary group $\{x_{i}|i\in[1,6]\}$ can be regarded as a problem statement similar to that in the RPM problem, while the test images and images in the auxiliary group $\{x_{i}|i\in[7,14]\}$ can be seen as an option pool anaLogous to the one in the RPM problem.
-This transformation process can be visualized through Figure 8.
+We regard Valen as a strong baseline for image progression pattern reasoning and as a model capable of tackling image clustering reasoning problem such as Bongard-Logo.
+To adapt Valen to Bongard-Logo, we transform the task accordingly.
+In a given Bongard-Logo instance, we denote the 14 images as $\{x_{i}|i\in[1,14]\}$.
+Specifically, the images in the primary group are $\{x_{i}|i\in[1,6]\}$, while the images in the auxiliary group are $\{x_{i}|i\in[8,13]\}$.
+Furthermore, $x_{7}$ is to be assigned to the primary group, and $x_{14}$ is to be assigned to the auxiliary group.
+This paper proposes that image clustering reasoning problems, such as Bongard-Logo, can be approached from different perspectives.
+Specifically, these tasks can be recast as option-evaluation problems, analogous to those in the RPM task.
+To elaborate, the images in the primary group $\{x_{i}|i\in[1,6]\}$ can be regarded as a problem statement akin to that in RPM, while the test images and those in the auxiliary group $\{x_{i}|i\in[7,14]\}$ act as an option pool analogous to the one in RPM.
+This transformation is shown in Figure 8.
 Figure 8: Transformation of Bongard-Logo problem
-After such transformation, Valen can be effectively applied to image clustering reasoning problems like the Bongard-Logo problem.
-It is noteworthy that when migrating Valen to Bongard-Logo, some adjustments are still necessary to accommodate the data characteristics of Bongard-Logo.
+After this transformation, Valen can be effectively applied to image clustering reasoning problems such as Bongard-Logo.
+It is noteworthy that when transferring Valen to Bongard-Logo, some adjustments are still necessary to accommodate the data characteristics of Bongard-Logo.
 Specifically, due to the scarcity of training instances in the Bongard-Logo problem, the ViT-based image representation extractor may struggle to perform effectively [21, 32].
@@
 Therefore, we have redesigned the image representation extractor, originally illustrated in Figure 4, into a convolutional network concatenated with a Transformer-Encoder.
-The new extractor design can be visualized in Figure 9.
+The new design is illustrated in Figure 9.
 Figure 9: Representation extraction of Bongard-Logo image
-In the figure 9, we can observe that when a Bongard-Logo image is processed by the new Valen’s feature extractor, it is first encoded into a feature map by ResNet18.
-Subsequently, this feature map is decomposed into multiple tokens and fed into a standard Transformer-Encoder to obtain multi-viewpoint representations.
-Such a design allows for the combination of CNN’s ability to analyze a smaller number of samples[1, 2, 3, 17] and ViT’s ability to analyze samples from multiple viewpoint[28, 24].
-This design has been applied in previous work[24].
+In this figure, when a Bongard-Logo image is processed by Valen’s new feature extractor, it is first encoded into a feature map by ResNet18.
+This map is then decomposed into multiple tokens and fed into a Transformer-Encoder to obtain multi-viewpoint representations.
+The design combines CNN’s sample efficiency with ViT’s multi-viewpoint modeling capacity [1, 2, 21, 32, 27] and has been used in previous work [27].
 Additionally, we have simplified the incomplete matrix progressive pattern extractor, which was originally intended to be structured as shown in Figure 6, into the form illustrated in Figure 10.
 Figure 10: Pattern extraction of Bongard-Logo incomplete matrix
-In the figure10, we can see that we have abandoned the use of MLP to extract the row and column progressive information from the incomplete matrix in Bongard-Logo, because the transformed Bongard-Logo problem only contains clustering patterns and not progressive patterns.
-Instead, we directly feed the representations of all images within the incomplete matrix of Bongard-Logo as tokens into the Transformer-Encoder to extract $M$ clustering patterns of Bongard-Logo.
-When Valen processes Bongard-Logo problems, $M$ is set to 6.
+In this figure, we abandon the use of MLP to extract the row and column progressive information from the incomplete matrix in Bongard-Logo, because the transformed Bongard-Logo problem only contains clustering patterns and not progressive patterns.
+We then feed the representations of all images within the Bongard-Logo incomplete matrix as tokens into the Transformer-Encoder to extract $M$ clustering patterns.
+When processing Bongard-Logo, $M$ is set to 6.
 V Tine, a Method for Advancing the Solution Distribution Boundary
-Current visual abstract reasoning problem solvers, including Valen, mostly follow a pattern: solvers are keen on assigning probabilities to each option of reasoning problems.
-The loss functions designed for these solvers are diverse, and the backbone network designs of these solvers vary greatly, but most of them adhere to this paradigm.
-We name this type of solver as “probability-highlighting solver”.
-Therefore, we can interpret these solvers as a conditional probability distributions obeyed by abstract reasoning problems, or understand them as providing distributions of solutions to abstract reasoning problems.
-If we use RPM solvers as examples, then these solvers can be abstracted as a conditional probability distribution, denoted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$.
-(Accordingly, when Valen is applied to Bongard-Logo problems, it can be abstracted as a conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{6};\theta)$).
-In this context, the two conditions of the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ are the statement of the reasoning problem $\{x_{i}\}_{i=1}^{8}$ and the optimizable parameters $\theta$ of the solver.
-Here, the $x_{\alpha}$ refers to the option that is pending verification.
-This conditional distribution implies that, when given a problem statement $\{x_{i}\}_{i=1}^{8}$ and an option $x_{\alpha}$, it is possible to calculate the probability of $x_{\alpha}$ being a solution to $\{x_{i}\}_{i=1}^{8}$.
-The optimization of probability-highlighting solver can be regarded as optimizing the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$.
-A wide range of loss functions can be used for optimization, but they all share the same objective: to assign high probabilities to the primary samples in the probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ and low probabilities to the auxiliary options.
+Current end-to-end visual abstract reasoning solvers, including Valen, are keen on assigning probabilities to each option.
+Although their loss functions are diverse and their backbone architectures vary significantly, most adhere to this paradigm.
+We term such solvers “probability-highlighting solvers”.
+Therefore, we can interpret these solvers as a conditional probability distribution governing the solutions to abstract reasoning problems.
+Taking RPM solvers as examples, they can be abstracted as a conditional distribution, denoted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$.
+(Accordingly, when Valen is applied to Bongard-Logo problems, it can be abstracted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{6};\theta)$).
+Here, $\{x_{i}\}_{i=1}^{8}$ is the statement of the reasoning problem, $\theta$ represents the learnable parameters of the solver, and $x_{\alpha}$ refers to the option that is pending verification.
+This distribution gives the probability that the solver’s parameters $\theta$ deem $x_{\alpha}$ to be the solution to the context $\{x_{i}\}_{i=1}^{8}$.
+The optimization of a probability-highlighting solver corresponds to tuning the parameters $\theta$ within the conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$.
+Although a wide range of loss functions can be employed, they all share the same objective: adjust parameters $\theta$ so that the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ assigns high probability to the primary samples and low probability to the auxiliary samples.
 Such an optimization objective makes us realize that the optimization process of the probability-highlighting solver is not aimed at fitting the distribution of correct solutions to visual abstract reasoning problems, but rather is dedicated to fitting a distribution jointly delimited by primary and auxiliary samples.
-To simplify the explanation of our findings, we assume that the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ follows a Gaussian distribution.
-Consequently, we can explicitly identify the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ as the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$.
-Then, the negative logarithm of the probability density function of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$ can be expressed as follows:
-$-\log(P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta}))=\frac{{(x_{a}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(|\,\sigma_{\theta}|\,)$ (1)
-Evidently, we have omitted the constant term $-1/2\log(2\pi)$.
-In this formula, $\mu_{\theta}$ and $\sigma_{\theta}$ respectively represent the mean and standard deviation of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$.
-Since we have defined the form of distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ as a Gaussian distribution, we consequently abstract the optimizable parameters $\theta$ into optimizable mean $\mu_{\theta}$ and standard deviation $\sigma_{\theta}$.
-Optimizing the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$ using the loss function is equivalent to optimizing the values of $\mu_{\theta}$ and $\sigma_{\theta}$.
-Our optimization objective is that when primary samples (the correct options in an RPM problem, or the samples within the main group in a Bongard-Logo problem) are plugged into formula (1), we expect to obtain a smaller value.
-Conversely, when auxiliary samples (the incorrect options in an RPM problem, or the samples within the auxiliary group in a Bongard-Logo problem) are plugged into formula (1), we hope that the formula returns a larger value.
-Evidently, constraining the return value of the formula to achieve a minimum is equivalent to optimizing the value of $\mu_{\theta}$ in the formula.
-Conversely, constraining the return value to achieve a maximum is equivalent to minimizing the value of $\sigma_{\theta}$ in the formula.
+To simplify the explanation of our findings, we assume that the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ follows a Gaussian distribution, so we write it explicitly as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$, where $\mu_{\theta}$, $\sigma_{\theta}$ represent the mean and standard deviation, respectively.
+Then, the negative logarithm of this Gaussian density can be expressed as follows:
+$-\log(P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta}))=\frac{{(x_{\alpha}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(2\pi|\,\sigma_{\theta}|\,)$ (1)
+Accordingly, the optimization objective is to tune the values of $\mu_{\theta}$ and $\sigma_{\theta}$ so that when primary samples (i.e., the correct RPM options or samples within the primary group in Bongard-Logo) are plugged into formula (1), the formula returns a smaller value; conversely, when auxiliary samples (i.e., the incorrect RPM options or the samples within the auxiliary group in a Bongard-Logo problem) are plugged into formula (1), the formula returns a larger value.
+Evidently, minimizing the return value of the formula (1) amounts to minimizing the Euclidean distance between $\mu_{\theta}$ and $x_{\alpha}$; maximizing it corresponds to minimizing the value of $\sigma_{\theta}$.
 In other words, the probability-highlighting solver positions the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ for the reasoning problem through primary samples, while the standard deviation of the solution distribution is constrained by auxiliary samples.
@@
 This discouraging reality can be visualized in Figure 11.
-Note that the figure 11 is intended to schematically represent our analysis results, rather than to depict the actual form of the solution distribution or the correct solution distribution.
+Note that this figure is intended to schematically represent our analysis results, rather than to depict the actual form of the solution distribution or the correct solution distribution.
 Figure 11: Schematic diagram of the relative relationship between the solution distribution, the correct solution distribution, and the primary and auxiliary samples.
 This paper argues that to further improve the reasoning accuracy of Valen in solving RPM problems and Bongard-Logo problems, it is necessary to provide Valen with more judicious auxiliary samples.
-By “more judicious auxiliary samples,” we refer to those that can better delimit the solution distribution provided by Valen.
+By “more judicious auxiliary samples”, we refer to those that can better delimit the solution distribution provided by Valen.
 These more judicious auxiliary samples have the potential to narrow down the standard deviation of the solution distribution.
@@
 Figure 12: Judicious auxiliary samples
-As seen in the figure 12, the judicious auxiliary samples we desire are those that appear more authentic or closer to the primary samples from Valen’s perspective, compared to the original auxiliary samples provided in the reasoning problems.
-Since the property of being “judicious” is evaluated from Valen’s perspective, whether an auxiliary sample is judicious or not, and the extent of its judiciousness, lacks interpretability from a human perspective.
-This implies that, relative to the primary samples, samples with smaller attribute shifts or fewer shifted attributes may not necessarily be considered more judicious from Valen’s perspective.
-Therefore, this paper argues that we can adopt an adversarial learning strategy to provide Valen with auxiliary samples that are more judicious from Valen’s perspective.
-Given this insight, we design Tine (Technique for edge Identification of solution distribution based on adversarial Networks) method.
-Specifically, we intend to emulate adversarial generation techniques similar to GAN[7] or WGAN[33] to generate more judicious samples for Valen.
-However, unlike GAN or WGAN, whose objective is to train robust and high-performance generators, our utilization of adversarial learning aims to refine Valen, which serves as the discriminator in this context.
+As shown in this figure, the judicious auxiliary samples we desire appear more authentic or closer to primary samples from Valen’s perspective than the original auxiliary samples.
+However, since judiciousness is evaluated from Valen’s perspective, it lacks human interpretability; samples with smaller or fewer attribute shifts are not necessarily deemed more judicious by Valen.
+To this end, this paper argues that we can use adversarial learning to supply Valen with auxiliary samples it deems more judicious.
+Given this insight, we design the Tine (Technique for edge Identification of solution distribution based on adversarial Networks) method.
+Specifically, we adopt adversarial generation techniques similar to GAN [9] or WGAN [39] to generate more judicious samples for Valen.
+However, unlike GAN or WGAN, whose objective is to train robust and high-performance generators, we use adversarial learning to refine Valen, which acts as the discriminator.
 Our objective is to design an auxiliary sample generator for Valen.
-The task of this generator is clear: to produce auxiliary samples that can achieve higher scores (probability) in Valen compared to the original auxiliary samples provided in the reasoning tasks.
-Conversely, Valen’s role is to classify these auxiliary samples as low-scoring ones.
-During the training process, the parameters of Valen and this generator are updated alternately and exclusively, with their performance gradually refined through adversarial learning, which is a common process in adversarial learning.
-Given that our pursuit is to generate more judicious auxiliary samples from Valen’s perspective, we are not concerned with their interpretability, attribute geometry, or whether they strictly conform to the nature of reasoning problems.
-Therefore, we believe that instead of forcing the generator to produce image-based auxiliary samples, which would incur greater computational overhead and optimization difficulties, it is more effective to generate representations of judicious auxiliary samples.
-By solely generating multi-viewpoint representations of reasoning problems, the generator’s structure becomes simpler and easier to optimize.
-Consequently, our generator is designed as shown in Figure 13.
-The figure 13 depicts Valen, which is tailored for addressing the RPM problem, and its accompanying auxiliary sample representation generator.
+Its task is to produce samples that receive higher probabilities from Valen than the original auxiliary samples.
+Accordingly, Valen’s role is to classify these generated samples as low-scoring ones.
+During training, the parameters of Valen and the generator are updated alternately and exclusively, undergoing standard adversarial refinement.
+Given that our goal is to generate more judicious auxiliary samples from Valen’s perspective, we are not concerned with their interpretability, attribute values, or whether they strictly conform to the nature of reasoning problems.
+Consequently, rather than forcing the generator to produce pixel-level auxiliary samples, which would incur greater computational overhead and optimization difficulties, we opt to generate representations of judicious auxiliary samples.
+By outputting only multi-viewpoint representations, the generator enjoys a simpler architecture and becomes easier to optimize.
+The generator is shown in Figure 13.
 Since we have already transformed the Bongard-Logo problem into an RPM-like problem, we will not further elaborate on the structure of the generator for the Bongard-Logo version.
 Figure 13: The structure of the generator associated with Valen
-As depicted in Figure 13, the left gray dashed box encloses the previously designed Valen, while the right gray dashed box encloses the generator designed for Valen.
-The figure 13 shows that the backbone of this generator is a standard Transformer-Encoder, which generates representations of more judicious auxiliary samples based on the representations of the abstract reasoning problem statement $\{r_{i}|i\in[1,8]\}_{v}$, the progressive patterns $\{P_{9m}|m\in[1,M]\}_{v}$ of the statement, and $L$ learnable vectors.
-Both the representations of the statement and their progressive patterns are provided by Valen, while the $L$ learnable vectors require additional establishment for the generator.
-It is noteworthy that when dealing with PGM problems in RPM tasks, we have designed special positional encoding for the representations of the problem statement.
+As depicted in Figure 13, the left gray dashed box encloses the Valen model, while the right gray dashed box encloses the generator designed for Valen.
+The backbone of this generator is a standard Transformer-Encoder, which generates representations of more judicious auxiliary samples based on the representations of the abstract reasoning problem statement $\{r_{i}|i\in[1,8]\}_{v}$, their progressive patterns $\{P_{9m}|m\in[1,M]\}_{v}$, and $L$ learnable vectors.
+Both the representations and the progressive patterns of the statement are provided by Valen, while the $L$ learnable vectors must be additionally introduced for the generator.
+It is noteworthy that when dealing with PGM problems within RPM tasks, we have designed special positional encoding for the representations of the problem statement.
 This positional encoding can be specifically attached to incomplete matrices with missing option positions and is set symmetrically along the diagonal.
-After the Transformer-Encoder processes the input representations (i.e., $\{r_{i}|i\in[1,8]\}_{v}$, $\{P_{9m}|m\in[1,M]\}_{v}$ and $L$ learnable vectors) and generates the output, we extract and save those $L$ optimizable vectors from the output.
-These $L$ (processed) output learnable vectors are then individually processed by an MLP into vectors of double length.
-These $L$ doubled-length vectors are transformed into $L$ vector using the reparameterization technique[8].
-The process of reparameterization, which is a practical approach used in VAE[8], can be visualized as Figure 14.
+After the Transformer-Encoder processes the input representations (i.e., $\{r_{i}|i\in[1,8]\}_{v}$, $\{P_{9m}|m\in[1,M]\}_{v}$ and $L$ learnable vectors) and generates the output, we extract and save those $L$ learnable vectors.
+These processed vectors then pass through an MLP that doubles their dimension.
+The resulting $L$ doubled-length vectors are transformed back into $L$ vectors using the reparameterization technique [10].
+The details of this technique are shown in Figure 14.
 Figure 14: Reparameterization
-These $L$ reparameterized vector are regarded as the representations of $L$ auxiliary samples.
+These $L$ reparameterized vectors are regarded as the representations of $L$ auxiliary samples.
 It is worth noting that $L$ is a hyperparameter that can be freely set, determining how many judicious auxiliary samples Tine generates for Valen in one instance.
-In this paper, $L$ is set to 9.
-In summary, our Tine method is proposed after summarizing the working principles of probability-highlighting solvers such as Valen.
-We believe that such solvers establish a solution distribution for each instance of a reasoning problem, with the optimization objective being to assign high probability to primary samples and low probability to auxiliary samples.
-This means that the optimization target of the solution distribution is not aligned with the distribution of correct solutions to the reasoning problems, but rather with the distribution delimited by both primary and auxiliary samples.
-Unfortunately, this delimited distribution is often much coarser than the correct solution distribution.
-To make the distribution delimited by both primary and auxiliary samples closer to the correct solution distribution, we designed Tine based on adversarial learning.
-Although subsequent experiments demonstrated that Tine can effectively improve Valen’s performance, we believe that the Tine method still fails to address the underlying essence of the problem.
-Therefore, further exploration is conducted in this paper.
+Here, $L$ is set to 9.
+In short, our Tine method was conceived after summarizing the working principles of probability-highlighting solvers.
+We contend that such solvers establish a solution distribution for every instance of a reasoning problem and adjust it so that primary samples receive high probability while auxiliary samples receive low probability.
+This means their optimization target is not to align with the distribution of correct solutions to the reasoning problems, but rather to align with the distribution jointly delimited by primary and auxiliary samples.
+Unfortunately, this delimited distribution is often much coarser than the correct one.
+To bring the former closer to the latter, we designed Tine based on adversarial learning.
 VI Funny: Method for Planning the Distribution of Reasoning Problem Samples.
@@
 The Tine method generates more judicious auxiliary samples to further constrain the standard deviation.
-Although experimental results demonstrate the effectiveness of Tine, it remains uncertain whether Tine constrains the standard deviation too little or not enough.
-Furthermore, Tine based on adversarial learning also encounters all the issues associated with adversarial learning, such as unstable training processes and difficulties in determining the completeness of training.
-Given this series of challenges, this paper proposes the Funny (Framework Utilizing Neural Networks for Yielding Image Representation Distribution) method.
-Our Funny method argues that instead of laboriously adopting adversarial learning strategies to make the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ provided by the solver closer to the correct solution distribution, it is more effective to directly specify the range of correct solutions or, alternatively, to explicitly define the details of the correct solution distribution.
-This paper will next analyze the definition of correct solutions for image progressive pattern reasoning problems such as RPM and image clustering reasoning problems such as Bongard-Logo.
-According to the design of RPM problems, the attributes of an image within the RPM problem are divided into key attributes related to reasoning and non-key attributes unrelated to reasoning.
-The design purpose of attributes unrelated to reasoning is to interfere with the solver’s reasoning and test its ability to filter attributes.
-This paper has discovered a point that cannot be neglected: attributes are the only factor that determines the specificity of RPM images.
-Any two RPM images with the same key attributes have the same identity in the reasoning process, regardless of how different their non-key attributes are.
-This implies that any sample whose key attributes pass the verification can be considered a correct solution to the RPM problem.
-For instance, in the RAVEN problem[15], the angle of an entity is considered a non-key attribute that is unrelated to reasoning, while attributes such as color, size, shape, and position are all key attributes.
-This leads to the situation where a white, medium-sized hexagon positioned in the center retains its reasoning identity regardless of how it is rotated.
-No matter how it rotates, it will not be transformed from a correct solution to an incorrect one or vice versa.
+However, it remains unclear whether this constraint is insufficient or excessive.
+Moreover, being rooted in adversarial learning, Tine inevitably inherits the associated drawbacks, such as unstable training and difficulty in assessing convergence.
+Faced with these challenges, we propose the Funny (Framework Utilizing Neural Networks for Yielding Image Representation Distribution) method.
+Our Funny method posits that, rather than laboriously adopting adversarial learning to make the solver’s distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ close to the correct solution distribution, it is more effective to directly specify the range of correct solutions or to explicitly characterize the correct solution distribution.
+Toward this goal, the present section dissects the correct solutions for both image progression pattern and image clustering reasoning problems.
+VI-A Key attributes and non-key attributes
+According to the design of RPM problems, the attributes of an image are divided into key attributes that are relevant to reasoning and non-key attributes that serve as distractors.
+The latter are introduced solely to interfere with the solver and to evaluate its ability to filter irrelevant information.
+It is critical to note that attributes alone determine the identity of an RPM image: any two images that share identical key attributes have the same identity in the reasoning process, regardless of their non-key differences.
+Consequently, any sample whose key attributes satisfy the governing pattern is deemed a correct solution.
+For example, in the RAVEN problem[16] the rotation angle of an entity is treated as a non-key attribute, whereas color, size, shape, and position are key; thus, a white, medium-sized hexagon positioned in the center retains its reasoning identity no matter how it is rotated and cannot be transformed by rotation from a correct solution into an incorrect one, or vice versa.
 The design of correct solutions for Bongard-Logo problems is much more complex.
-The reasoning identity of a Bongard-Logo sample is determined by the context within the specific instance it belongs to.
-Among the various attributes of a sample, which are key attributes and which are non-key attributes are influenced by the context within the relevant instance.
-In Bongard-Logo problems, some instances focus on more macroscopic attributes of the image, such as its closure and concavity & convexity; while other instances focus on more detailed attributes, such as the number of entities in the image and the order of their arrangement.
-This is different from RPM problems, where the key and non-key attributes are consistent across all instances within the RPM.
-Given the differences in determining correct solutions, or in identifying key and non-key attributes, between Bongard-Logo and RPM problems, the application of our proposed Funny approach differs when applied to these two types of problems.
-We will discuss them separately.
-VI-A Funny for RPM problems
+The reasoning identity of a Bongard-Logo sample is determined by its instance context, which also dictates the key and non-key attributes.
+Some instances focus on macroscopic attributes such as closure, concavity and convexity, while others concentrate on detailed attributes such as the number of entities and their spatial order.
+This contrasts with RPM problems, where key and non-key attributes are consistent across all instances.
+Due to the differences between Bongard-Logo and RPM problems in determining correct solutions and in distinguishing key from non-key attributes, we apply tailored variants of our Funny approach to the two problem types; each will be discussed separately.
+VI-B Funny for RPM problems
 Given that the key attributes and non-key attributes in the RPM problem are fixed, with key attributes being key across all instances and non-key attributes being unfavored in any instance, this paper proposes that the representations of samples with the same key attributes can be mapped into a single Gaussian distribution.
 Within this Gaussian distribution, the key attributes of the samples are identical, but the non-key attributes vary.
-Consequently, the distribution of representations for all samples in the entire RPM problem is formulated as a mixture of Gaussian distributions.
-By exposing Valen to this mixture of Gaussian distributions instead of individual sample representations, we effectively clarify the distribution of correct solutions for Valen.
-Our intention is to achieve the transformation of the sample representation space as illustrated in the Figure 15.
+Consequently, the distribution of representations for all samples in the entire RPM problem is formulated as a mixture of Gaussian distributions [45].
+By exposing Valen to this mixture instead of individual sample representations, we effectively clarify the distribution of correct solutions for Valen.
+Our intention is to achieve the transformation of the sample representation space as illustrated in Figure 15.
 Figure 15: Modeling the representation distribution as a mixture of Gaussian distributions.
-Therefore, the purpose of the Funny method is to plan the distribution of samples in the RPM problem as a mixture of Gaussian distributions, with the key attributes serving as the component means.
-The design of Valen’s Funny method, applied to tackle the RPM problem, is quite straightforward.
-The Funny method separately extracts the key attributes and non-key attributes of RPM images, treating the key attributes as the center and the non-key attributes as biases relative to this center.
-Therefore, we refer to the representation of the key attributes extracted from the RPM image $x_{i}$ as $center_{i}$, and the representation of the non-key attributes as $bias_{i}$.
-Interestingly, the sum of $center_{i}$ and $bias_{i}$ can be viewed as samples drawn from a Gaussian distribution $N(center_{i},\sigma^{2})$, which represents a distribution with a mean of $center_{i}$ and an unknown variance $\sigma^{2}$.
+Therefore, the Funny method aims to plan the RPM sample distribution as a Gaussian mixture in which the key attributes define the component means.
+The design of the Funny method for tackling RPM problems is straightforward: Funny adopts the image attribute extraction scheme illustrated in Figure 16.
+Figure 16: The form of image attribute extraction adopted by Funny.
+In this figure, the Funny method extracts the key attributes and non-key attributes of RPM images, treating the key attributes as the center and the non-key attributes as biases relative to this center.
+Accordingly, we refer to the representation of the key attributes extracted from the RPM image $x_{i}$ as $center_{i}$, and that of the non-key attributes as $bias_{i}$.
+Interestingly, the sum of $center_{i}$ and $bias_{i}$ can be viewed as a sample drawn from the Gaussian distribution $N(center_{i},\sigma^{2})$, which represents a distribution with a mean of $center_{i}$ and an unknown variance $\sigma^{2}$.
 When the $bias_{i}$ follow a standard normal distribution, this Gaussian distribution can be considered as $N(center_{i},1)$, which is a distribution with a mean of $center_{i}$ and a variance of 1.
-In this scenario, representations of samples with the same key attributes but different non-key attributes only differ in their biases.
-This implies that the Funny method successfully maps representations of samples with the same key attributes into the same Gaussian distribution.
+In this scenario, representations of samples with the same key attributes but different non-key attributes only differ in their biases, implying that the Funny method successfully maps representations of samples with the same key attributes into the same Gaussian distribution.
 Based on this hypothesis, we design a multi-viewpoint image representation extractor as illustrated in Figure 17, which replaces Valen’s original representation extractor when equipped with the Funny method.
 Figure 17: Funny representation extractor
-As can be seen in the figure 16, Funny encodes the image sample $x_{i}$ of the reasoning problem into two sets of homomorphic multi-viewpoint representations, named $center_{i}$ and $bias_{i}$ respectively.
-The final multi-viewpoint representation $r_{i}$ of the image $x_{i}$ extracted by the Funny representation extractor is obtained by summing $center_{i}$ and $bias_{i}$.
-However, the new representation extractor of the Funny method, which encodes images $x_{i}$ into $center_{i}$ and $bias_{i}$, still faces two challenges: 1.
-How to ensure that the $center_{i}$ stores key attributes while the $bias_{i}$ stores non-key attributes?
-2.
-Given that the union of key and non-key attributes constitutes all attributes, how to guarantee that the combination of $center_{i}$ and $bias_{i}$ can store all attributes of the image $x_{i}$?
-These two questions pose challenges to both the center and bias individually and collectively.
+As can be seen in this figure, Funny encodes the reasoning image $x_{i}$ into two sets of homomorphic multi-viewpoint representations, termed $center_{i}$ and $bias_{i}$ respectively.
+The final multi-viewpoint representation $r_{i}$ is obtained by element-wise summation of $center_{i}$ and $bias_{i}$.
+However, the new representation extractor of the Funny method, which encodes images $x_{i}$ into $center_{i}$ and $bias_{i}$, still faces two challenges: (1) How to ensure that the $center_{i}$ stores key attributes while the $bias_{i}$ captures non-key attributes?
+(2) Given that the union of key and non-key attributes constitutes all attributes, how to guarantee that the combination of $center_{i}$ and $bias_{i}$ can store all attributes of the image $x_{i}$?
+These questions pose challenges to both the center and bias individually and collectively.
 This paper argues that by enforcing $bias_{i}$ to follow a standard normal distribution $N(0,1)$ and simultaneously enforcing all random sampling results from a Gaussian distribution $N(center_{i},1)$ to assume the same reasoning identity, it can be ensured that $center_{i}$ stores at least the key attributes.
-This is because, when samples drawn from Gaussian distribution $N(center_{i},1)$ all share the same reasoning identity, $bias_{i}$, which follows a standard normal distribution $N(0,1)$, ceases to influence the reasoning, and $center_{i}$ takes full responsibility for the reasoning task, indicating that $center_{i}$ stores at least the key attributes at this point.
-Furthermore, to ensure that the combination of $center_{i}$ and $bias_{i}$ can store all attributes within image $x_{i}$, and consequently guarantee that the $bias_{i}$ stores the image attributes omitted by the $center_{i}$, it suffices to maximize the mutual information between combination of $center_{i}$ and $bias_{i}$ and all attributes within image $x_{i}$.
-The most direct method to maximize the mutual information between image attributes and image representations is to append an additional reconstruction task.
-Therefore, when the Valen framework incorporates the Funny method, its structure should be as illustrated in the Figure 17.
+Under this constraint, $bias_{i}$, which follows a standard normal distribution $N(0,1)$, ceases to influence the reasoning, and $center_{i}$ takes full responsibility for the reasoning task, indicating that $center_{i}$ stores at least the key attributes at this point.
+Furthermore, we propose that maximizing the mutual information between $center_{i}+bias_{i}$ and the image $x_{i}$ suffices to ensure that this combined representation retains the complete set of attributes within $x_{i}$, thereby guaranteeing that $bias_{i}$ encodes the attributes omitted by $center_{i}$.
+The most direct method to maximize this mutual information is to add a reconstruction task.
+Therefore, when the Valen framework incorporates the Funny method, the overall architecture is as shown in Figure 18.
 Figure 18: The Valen framework incorporates the Funny method
-In the figure 17, we can observe that the representation of key attribute, which is $\{center_{i}|i\in[1,9]\}$, extracted by the Funny extractor for a $3\times 3$ image matrix $\{x_{i}|i\in[1,9]\}$ is superimposed with the sampling results $\epsilon$ from a Gaussian normal distribution $N(0,1)$ and then fed into the regular Valen framework to undertake subsequent reasoning tasks.
-This is done to clarify with the help of Valen that the representations drawn from the Gaussian distribution $N(center,1)$ have the same reasoning identity.
-Furthermore, in the figure 17, we can observe that Funny introduces three additional regression tasks based on mean squared error (MSE) loss to Valen, which are indicated by dashed lines in red, yellow, and green.
-The green dashed line indicates the regression task between the representation of non-key attributes, which is $\{bias_{i}|i\in[1,9]\}$, and the random sampling results $\epsilon$ from the standard normal distribution $N(0,1)$.
-It is worth noting that the $\epsilon$ here needs to be resampled and cannot be the one previously superimposed on the $\{center_{i}|i\in[1,9]\}$.
-The objective of this green task is to enforce the bias to conform to the standard normal distribution $N(0,1)$.
-The red and yellow dashed lines indicate the two regression tasks performed by the Funny decoder.
-The task indicated by the red dashed line aims to reconstruct the corresponding RPM image $x_{i}$ from the representation obtained by adding $center_{i}$ and $bias_{i}$.
-The task indicated by the yellow dashed line aims to reconstruct RPM images from the representation obtained by adding $center_{i}$ and $\epsilon$.
+As shown in this figure, the key attribute representations $\{{center}_{i}|i\in[1,9]\}$ extracted from the $\times3$ image matrix $\{x_{i}|i\in[1,9]\}$ are superimposed with $\epsilon\sim N(0,1)$ and then fed into the Valen framework for subsequent reasoning.
+This process uses the Valen model to confirm that representations sampled from $N(center_{i},1)$ share the same reasoning identity.
+Moreover, Funny introduces three additional regression tasks based on mean squared error (MSE) loss to Valen; these tasks are indicated by dashed lines in red, yellow, and green.
+The green dashed line indicates the regression task between the representation of non-key attributes $\{bias_{i}|i\in[1,9]\}$ and fresh random samples $\epsilon\sim N(0,1)$.
+The objective of this green task is to enforce $bias_{i}$ to follow $N(0,1)$.
+The red and yellow dashed lines indicate two regression tasks performed by the Funny decoder.
+The red task reconstructs the original RPM image $x_{i}$ from the combined representation $center_{i}+bias_{i}$, while the yellow task reconstructs $x_{i}$ from $center_{i}+\epsilon$.
 The objectives of these two reconstruction tasks are vastly different.
-The reconstruction task indicated by the red dashed line is aimed at maximizing the conditional distribution $P(x_{i}|center_{i}+bias_{i})$ during the training process of the Funny decoder.
-On the other hand, the reconstruction task indicated by the yellow dashed line is intended to ensure that the output of the Funny decoder follows the distribution of RPM images, thereby guaranteeing that the Funny decoder can effectively function as a distribution transformation function.
-The objective of the red reconstruction task, which is to maximize the conditional distribution $P(x_{i}|center_{i}+bias_{i})$, is equivalent to maximizing the mutual information between the attributes of $x_{i}$ and $center_{i}+bias_{i}$.
+The red task maximizes the conditional distribution $P(x_{i}|center_{i}+bias_{i})$ during training, which is equivalent to maximizing the mutual information between the attributes of $x_{i}$ and $center_{i}+bias_{i}$.
 This mutual information can be calculated using the following formula:
@@
 where $H(x_{i})$ is the entropy of $x_{i}$, and $H(x_{i}|center_{i}+bias_{i})$ is the conditional entropy of $x_{i}$ given $center_{i}+bias_{i}$.
-This formula quantifies the amount of information shared between $x_{i}$ and the combination of ${center}_{i}$ and ${bias}_{i}$.
-Additionally, the operations similar to the yellow reconstruction task, which involve constraining the output of the decoder to follow a certain distribution through MSE-based reconstruction tasks, have already been well-established and widely used in VAE[8].
-Our design intention for Funny is quite clear.
-We hope that, with the assistance of Valen, Funny can compel $center_{i}$ to store more key attributes related to reasoning, while $bias_{i}$ accumulates more non-key attributes unrelated to reasoning that could potentially interfere with it.
-By doing so, we aim to model the distribution of multi-viewpoint representations, generated by the $center_{i}+bias_{i}$, as a mixed Gaussian distribution that aids Valen in its reasoning process.
-The relationship between Funny and Valen is mutually beneficial and symbiotic.
+The yellow task forces the output of the Funny decoder to follow the distribution of RPM image pixels, ensuring that the Funny decoder acts as an effective distribution transformation function.
+Such MSE-based distribution matching is widely adopted in VAEs [10].
+Our design intention for Funny is clear: with the assistance of Valen, Funny compels $center_{i}$ to store more key attributes related to reasoning, while $bias_{i}$ accumulates more non-key attributes that could cause interference.
+Subsequently, the distribution of multi-viewpoint representations produced by $center_{i}+bias_{i}$ is planned as a mixture of Gaussians, thereby feeding back to Valen and enhancing its reasoning.
+The relationship between Funny and Valen is symbiotic.
 Although subsequent experiments demonstrated that Funny is a significantly effective method, given the fact that obtaining disentangled representations under unsupervised learning is not achievable [42], this paper holds a pessimistic attitude towards the extent of Funny’s decoupling of key and non-key attributes.
-VI-B Funny not funny
+VI-C Funny not funny
 What is of paramount importance is that, a Funny decoder with a conventional structure cannot bear the aforementioned theoretical framework.
-If we apply a deconvolution structure to the backbone of the Funny decoder, its optimization becomes extremely challenging, making it difficult to achieve satisfactory regression results and supervise the mutual information between $center_{i}+bias_{i}$ and $x_{i}$.
-Alternatively, if we replace the Funny decoder with a ViT structure symmetric to the representation extractor, although we obtain excellent regression results, it leads to a decrease in the normal reasoning accuracy of Valen instead of an increase.
-Unfortunately and unfunny, there is a gep between theory and practice.
-This paper speculates that this may be due to the impact of the regression task in Funny on the parameter optimization of the reasoning task in Valen.
-If we choose the normal ViT as the Funny decoder, the introduced regression task will be completed before the Valen’s reasoning task, because the optimization speed of the regression task is faster than that of the reasoning task at this time.
-We posit that the premature completion of the regression task may result in the $center_{i}$ storing more non-key attribute.
-Because we have set up the regression task from $center_{i}+\epsilon$ to $x_{i}$.
-This paper speculates that this might be the issue: when the regression task approaches completion prematurely, it may lead to $center_{i}$ falling into the gradient vanishing zone, subsequently rendering it difficult for $center_{i}$ to make further optimizations for the reasoning task.
-The non-key attribute stored in $center_{i}$ can disrupt the reasoning process of Valen.
+If we apply a deconvolution structure to the backbone of the Funny decoder, its optimization becomes extremely challenging, and thus hinders satisfactory regression and mutual-information supervision between $center_{i}+bias_{i}$ and $x_{i}$.
+Alternatively, if we replace the Funny decoder with a ViT structure symmetric to the representation extractor, we obtain excellent regression, yet Valen’s normal reasoning accuracy decreases rather than increases.
+Unfortunately and “unfunny”, there remains a gap between theory and practice.
+We speculate that this may arise from the regression task in Funny interfering with the parameter optimization of the reasoning task in Valen.
+If we adopt the standard ViT as the Funny decoder, the optimization speed of the regression task would be relatively fast, which could lead to the introduced regression task nearing completion before the reasoning task of Valen.
+We argue that this premature convergence is the crux: when the regression task approaches completion prematurely, it may lead to $center_{i}$ falling into the gradient vanishing zone, subsequently rendering it difficult for $center_{i}$ to make further optimizations for the reasoning task; the non-key attribute stored in $center_{i}$ can disrupt the reasoning process of Valen.
 An attempt was made to linearly search for the scaling coefficient between the MSE-based regression task introduced by Funny and the CE-based reasoning task in Valen, but it did not yield an effective solution.
@@
 Fortunately and funny, this paper presents a novel decoder architecture, which we term the “Half-split Decoder” to address this issue.
-The structure of our designed Half-split Decoder is illustrated in the Figure 18.
+The structure of the Half-split Decoder is illustrated in the Figure 19.
 Figure 19: Half-split Decoder
-The upper half, separated by the blue dashed line in figure 18, represents the representation extractor of Funny, which has been previously shown in Figure 16.
-The lower half depicts the structure of our designed Half-split Decoder.
-In the figure 18, we can observe that our Half-split Decoder is composed of two identical structures, with each structure’s parameters optimized independently without parameter sharing.
-It is evident that each individual structure resembles an inverted version of a ViT[17], which we will not delve into further in this paper.
-These two identical structures regress the multi-viewpoint representation $r_{i}$ into two outcomes, $xu_{i}$ and $xd_{i}$, respectively.
-Here $r_{i}$ can be derived either by summing ${center}_{i}$ with $\epsilon$ or by summing ${center}_{i}$ and $bias_{i}$, corresponding to the two regression requirements of Funny.
-However, the core of the Half-split Decoder lies not in these structures but in the two distinct activation functions illustrated in the figure 18.
-We represent these two activation functions as follows:
+The upper half, separated by the blue dashed line, shows Funny’s representation extractor, previously illustrated in Figure 17. The lower half depicts the Half-split Decoder.
+As shown in Figure 19, the Half-split Decoder is composed of two identical structures, with each structure’s parameters optimized independently without parameter sharing.
+Each branch resembles an inverted version of a ViT[21], which we will not discuss further.
+The two branches regress the multi-viewpoint representation $r_{i}$ into two outcomes, $xu_{i}$ and $xd_{i}$, respectively.
+Here $r_{i}$ can be derived either by summing ${center}_{i}$ with $\epsilon$ or by summing ${center}_{i}$ and $bias_{i}$, corresponding to the two regression targets of Funny.
+The core of the Half-split Decoder, however, lies not in these branches but in the two distinct activation functions depicted in Figure 19.
+We express these two activation functions as follows:
 $f_{1}(xu_{i})=\frac{\text{Sigmoid}(xu_{i})+1}{2},f_{2}(xd_{i})=\frac{\text{Sigmoid}(xd_{i})}{2}$ (3)
-In the above equation, it is evident that the value range of $f_{1}(xu_{i})$ lies within $(0.5,1)$, while the value range of $f_{2}(xd_{i})$ belongs to $(0,0.5)$.
-Although the value ranges of both $f_{1}(xu_{i})$ and $f_{2}(xd_{i})$ differ from the value range of the original RPM image $x_{i}$, which is $(0,1)$, we still compute the MSE loss between $f_{1}(xu_{i})$ and $x_{i}$, as well as between $f_{2}(xd_{i})$ and $x_{i}$.
+Clearly, $f_{1}(xu_{i})$ ranges over $(0.5,1)$, while $f_{2}(xd_{i})$ ranges over $(0,0.5)$.
+Although both ranges differ from the original RPM image pixel range $(0,1)$, we still compute the MSE loss between $f_{1}(xu_{i})$ and $x_{i}$, as well as between $f_{2}(xd_{i})$ and $x_{i}$.
 The sum of these two errors is used as the new loss function for the two regression tasks associated with the Funny decoder.
-This loss function can be represented by the following equation:
+This loss can be expressed as:
 $\ell_{\text{Funny}}=\text{MSE}(f_{1}(xu_{i}),x_{i})+\text{MSE}(f_{2}(xd_{i}),x_{i})$ (4)
-When both $xu_{i}$ and $xd_{i}$ in the formula are decoded from ${center}_{i}+{bias}_{i}$, the loss function is used to supervise the Mutual Information $\text{MI}(x_{i},{center}_{i}+{bias}_{i})$.
-Conversely, when both $xu_{i}$ and $xd_{i}$ are decoded from ${center}_{i}+\epsilon$, the loss function is employed to supervise the distribution that the output of the Funny decoder follows.
-This Half-split Decoder can ensure that the loss of the regression task introduced by Funny does not decrease to zero in the early stage of the reasoning task of Valen, thereby preventing the multi-view representation from becoming stagnant.
-VI-C Funny for Bongard-Logo problem
-Funny is inherently a method that analyzes the key degree of sample attributes and modifies the sample representation space accordingly.
-Since the key degree of an attribute in Bongard-Logo is determined by contextual information, the previous form of Funny is difficult to directly transfer to the Bongard-Logo problem.
-Therefore, we innovate the Funny method.
-This paper argues that we can seek a series of unlearnable Bongard-Logo image sample mapping functions $F(x_{i})$ that do not alter the key degree of image attributes in the current contextual environment.
-Additionally, these functions do not change the key attributes of the input image but produce varying degrees of alteration to non-key attributes.
-Clearly, we have not abandoned the motivation of Funny: to make it explicit to Valen that key attributes determine reasoning identity, while non-key attributes do not affect reasoning identity.
-The previous form of Funny achieved this by mapping the representations of samples with the same key attributes into a Gaussian distribution with a variance of 1.
+When both $xu_{i}$ and $xd_{i}$ are decoded from ${center}_{i}+{bias}_{i}$, $\ell_{\text{Funny}}$ supervises the Mutual Information $\text{MI}(x_{i},{center}_{i}+{bias}_{i})$, corresponding to Funny’s red regression task; when both $xu_{i}$ and $xd_{i}$ are decoded from ${center}_{i}+\epsilon$, $\ell_{\text{Funny}}$ supervises the distribution that the output of the Funny decoder follows, corresponding to the yellow regression task.
+The mismatch among the value ranges of $f_{1}(xu_{i})$, $f_{2}(xd_{i})$, and $x_{i}$ ensures that Funny’s regression loss $\ell_{\text{Funny}}$ does not drop to zero during the early stages of Valen’s reasoning, thereby preventing the multi-view representation $r_{i}$ from becoming stagnant.
+VI-D Funny for Bongard-Logo problem
+Funny is essentially a method that modifies the sample representation space based on the key status of sample attributes.
+Since this key status is context-dependent in Bongard-Logo, the original Funny cannot be directly transferred to this task.
+To this end, we argue that identifying a family of unlearnable Bongard-Logo image sample mapping functions $F(x_{i})$ that preserve the key attributes under the current context while introducing controlled perturbations to non-key attributes effectively extends Funny’s methodology to this problem.
+Clearly, we retain Funny’s original principle: to explicitly signal to Valen that key attributes determine reasoning identity, while non-key attributes do not affect it.
 So, how do we find such functions?
-Through Human-in-the-Loop (HITL).
-HITL refers to situations where human experts or users intervene and provide feedback or guidance during the operation or training of machine learning or deep learning models.
-By analyzing the three coexisting concepts in Bongard-Logo, we subsequently discovered that some direct data augmentation methods, such as horizontal or vertical flipping, and 90-degree, 180-degree, and 270-degree rotations, can effectively serve as the function $F(x_{i})$.
-Therefore, when we intend to apply the Funny method to Valen for addressing the Bongard-Logo problem, it suffices to randomly apply the aforementioned data augmentation methods to the training samples $x_{i}$ during the regular training process of Valen.
-It is noteworthy that, unlike in the Bongard-Logo problem, these data augmentation methods may disrupt the key attributes in RPM problem.
-In other words, the Funny method tailored for the Bongard-Logo problem cannot be transferred to RPM problems, and vice versa.
+Through Human-in-the-Loop (HITL)[43].
+HITL refers to situations where human experts or users intervene and provide feedback or guidance during the operation or training of machine learning or deep learning models[43, 44].
+By analyzing the three coexisting concepts in Bongard-Logo, we found that some straightforward augmentation techniques, such as horizontal or vertical flipping, and 90-degree, 180-degree, and 270-degree rotations, can effectively serve as the function $F(x_{i})$.
+Subsequently, to apply Funny to Valen on Bongard-Logo, we simply need to apply these augmentations randomly to the training samples $x_{i}$ during standard Valen training.
+It should be noted that, unlike in Bongard-Logo, the same augmentations can corrupt key attributes in the RPM problem; hence the Funny variant tailored for Bongard-Logo is not transferable to RPM, and vice versa.
 VII SBR: Method for Planning the Distribution of Reasoning Pattens.
-This paper suggests that we should also plan the distribution of the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ extracted by Valen for visual abstract reasoning problems, so that it is closer to the distribution of the correct solution’s pattern.
-However, there are no key or non-key parts in the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ encoded by Valen for RPM problems or Bongard-Logo problems, and even currently, the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ lacks any interpretability.
-Therefore, our Funny method is difficult to directly apply to the pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$.
-Therefore, we have designed a new method, SBR (Supervised Representation Distribution Planning Method).
-This method can explicitly model the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$ as a mixture of Gaussian distributions.
-The idea of SBR is straightforward: SBR defines the distribution form of $\{P_{nm}|m\in[1,M]\}_{v}$ as a mixture of Gaussian distributions and provides additional supervisory signals for the mean and variance of each sub-component in this hypothetical distribution.
-When establishing a Gaussian Mixture Model (GMM), providing means, covariance matrices, and component weights is a necessary procedure.
-However, the objective of this paper is to model the distribution of pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$ as a more interpretable and orderly mixture of Gaussian distributions, which involves altering the original form of the $\{P_{nm}|m\in[1,M]\}_{v}$ distribution.
-This implies that we need to provide a reasonable number of components for the hypothesized distribution form of $\{P_{nm}|m\in[1,M]\}_{v}$, namely the mixture of Gaussian distributions, assign reasonable means and variances to these components, and provide the covariance matrices between the components.
-Otherwise, forced and incorrect distribution modeling approaches may detrimentally affect the accuracy of Valen’s reasoning.
-It is noteworthy that in the RPM problems, each instance is not only designed with problem statements, option pools, and correct option labels, but also accompanied by metadata.
-This metadata provides a detailed description of the progressive patterns inherent in each instance.
-Interestingly, the intended meaning of $\{P_{nm}|m\in[1,M]\}_{v}$ encoded by Valen for an RPM instance also corresponds to a progressive pattern.
-Therefore, we decide to use metadata to provide relevant information about the target form of the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$.
-By enumerating all the metadata within the instances of the RPM problem, the enumerated metadata list can offer ideal information about the component centers for the target distribution form of $\{P_{nm}|m\in[1,M]\}_{v}$.
-These enumeration results are mutually independent, owing to the characteristics of metadata.
-In short, we enumerate all the metadata within the instances of the RPM problem to obtain a metadata list.
-After processing each element in the list as a Gaussian distribution, the entire list forms a mixture of Gaussian distribution.
-Subsequently, we naturally assign $\{P_{nm}|m\in[1,M]\}_{v}$ to the corresponding component of this mixture model, thereby achieving our goal of planning the distribution of pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$.
-This process can be visualized as Figure 19.
+This paper suggests that we should also plan the distribution of the pattern representation $\{P_{nm}|m\in[1,M]\}_{v}$ extracted by Valen for visual abstract reasoning problems, so that it aligns with the distribution of the correct solution’s pattern.
+However, there are no discernible key or non-key attributes in this pattern representation, and it still lacks sufficient interpretability.
+Therefore, our Funny method is difficult to directly apply to this planning task
+To this end, we have designed a new method, SBR (Supervised Representation Distribution Planning Method).
+This method can explicitly model the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$ as a mixture of Gaussian distributions[45].
+Its idea is straightforward: we reshape the original distribution of $\{P_{nm}|m\in[1,M]\}_{v}$ into an interpretable, well-ordered Gaussian mixture and explicitly supervise the mean and variance of each sub-component in this reshaped distribution.
+To establish a mixture of Gaussian distributions, it is necessary to provide means, covariance matrices, and component weights [46].
+Consequently, the SBR method needs to specify a reasonable number of components for the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$, assign appropriate means and variances to these components, and specify the covariance structure.
+Otherwise, any forced or incorrect specification of the reshaped distribution can degrade Valen’s reasoning performance.
+It is noteworthy that in RPM problems, each instance includes not only the problem statement, option pool, and index of correct option, but also metadata that provide a detailed description of the progressive patterns inherent in the instance.
+Interestingly, the intended semantics of $\{P_{nm}|m\in[1,M]\}_{v}$ encoded by Valen for an RPM instance also corresponds to these progressive patterns.
+Therefore, we choose to utilize the metadata to provide detailed information about the target form of the distribution of $\{P_{nm}|m\in[1,M]\}_{v}$.
+Specifically, we enumerate all the metadata within RPM instances to obtain a metadata list.
+After processing each element in the list as a Gaussian distribution, the entire list forms a Gaussian mixture.
+We then assign the pattern representations $\{P_{nm}|m\in[1,M]\}_{v}$ to the corresponding component of this mixture, thereby achieving our goal of planning the distribution of pattern representations.
+The assigning process can be visualized as Figure 20.
 Figure 20: The framework of SBR
-In the figure 19, it can be seen that we assume the number of elements in the metadata list to be $S$.
-Since there are often multiple decoupled progressive patterns within the RPM matrix, such as the “shape” and “line” progressive patterns present in the PGM problem[16], and the related metadata fully describes these progressive patterns, we decide to dissect all the progressive patterns recorded in the metadata and use them to individually guide the assignment of the $M$ progressive patterns stored in the $\{P_{nm}|m\in[1,M]\}_{v}$ to their respective distribution components.
-In the figure, we dissect the two progressive patterns, “pattern 3” and “pattern S”, recorded in the metadata and use them separately to guide the distribution attribution of $\{P_{n1}|n\in[1,9]\}_{v}$ and $\{P_{n2}|n\in[1,9]\}_{v}$.
-Evidently, when Valen is combined with the SBR method, the value of $M$ is no longer an arbitrarily set hyperparameter, but rather needs to be determined based on the types of decoupled progressive patterns stored in the metadata.
-For example, when the SBR method is applied to the PGM problem, the value of $M$ in Valen needs to be set to 2.
-Regarding how we process each element in the enumerated metadata list into distributions, we choose to use a standard Transformer-Encoder to encode the elements, which are descriptions of the progressive patterns, into vectors directly.
-A list containing $S$ elements will be encoded into $S$ vectors, which we denote as $\{q_{s}|s\in[1,S]\}$ here.
-If the first half of $q_{s}$ is regarded as the mean of a Gaussian distribution and the second half as the logarithm of variance, then $\{q_{s}|s\in[1,S]\}$ can represent a mixture of Gaussian distributions.
-We can again utilize the reparameterization technique shown in Figure 14 to reparameterize $\{q_{s}|s\in[1,S]\}$ into a new set of vectors, which we denote as $\{Q_{st}|s\in[1,S],t\in[1,T]\}$.
-In $\{Q_{st}|s\in[1,S],t\in[1,T]\}$, it can be observed that we reparameterized $\{q_{s}|s\in[1,S]\}$ $T$ times, where $T$ is set to 10 in this paper.
+As shown in this figure, we denote the number of elements in the metadata list to be $S$.
+Since there are often multiple decoupled progressive patterns within the RPM instance, such as the “shape” and “line” patterns present in the PGM task, and the accompany metadata fully describe these patterns, we decompose all patterns recorded in the metadata and use them individually to guide the assignment of the $M$ patterns stored in the $\{P_{nm}|m\in[1,M]\}_{v}$ to their respective distribution components.
+Accordingly, we decompose the two progressive patterns, “pattern 3” and “pattern SS”, recorded in the metadata and use them separately to guide the distribution attribution of $\{P_{n1}|n\in[1,9]\}_{v}$ and $\{P_{n2}|n\in[1,9]\}_{v}$.
+Evidently, when Valen is combined with the SBR method, the value of $M$ is no longer an arbitrary hyperparameter, but must be set to the number of decoupled progressive patterns stored in the metadata; for example, $M=2$ when the SBR method is applied to the PGM task.
+Regarding how we process each element in the enumerated metadata list into distributions, we employ a standard Transformer-Encoder to map the elements, which are descriptions of the progressive patterns, into vectors directly.
+A list of $S$ elements thus yields $S$ vectors, denoted $\{q_{s}|s\in[1,S]\}$.
+By treating the first half of each $q_{s}$ as the mean and the second half as the logarithm of variance, $\{q_{s}|s\in[1,S]\}$ defines a Gaussian mixture distribution.
+We again use the reparameterization technique shown in Figure 14 to reparameterize $\{q_{s}|s\in[1,S]\}$ $T$ times into a new set of vectors, denoted as $\{Q_{st}|s\in[1,S],t\in[1,T]\}$, with $T$ is set to 10.
 Regarding the process indicated by the light blue and purple lines in the Figure 20, where $\{P_{n1}|n\in[1,9]\}_{v}$ and $\{P_{n2}|n\in[1,9]\}_{v}$ are separately compressed into the corresponding Gaussian distribution $q_{3}$ and $q_{S}$, we achieve this by setting an additional loss function.
@@
 ${\ell_{\text{SBR}}}$ $=-\sum_{m=1}^{M}\sum_{\tilde{s}=1}^{S}\sum_{n=1}^{9}\sum_{t=1}^{T}\text{meta}_{{m\tilde{s}}}\cdot\log\frac{{{e^{({\bar{P}_{nm}}\cdot{Q_{\tilde{s}t}})/\tau}}}}{{\sum\nolimits_{s=1}^{S}{{e^{({\bar{P}_{nm}}\cdot{Q_{{s}t}})/\tau}}}}}$ (6)
-Where $\text{meta}_{{m\tilde{s}}}$ represents whether the the $m$-th progressive pattern recorded in the metadata of current instance matches the $\tilde{s}$-th element in the metadata list and $\tau$ is a learnable parameter.
-As can be seen from the formula (5), considering the overhead of computational resources, we do not individually compress each viewpoint in $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$ into the constructed mixture of Gaussian distributions, but instead first calculate the viewpoint average $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$ of $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}$ and then plan the distribution of $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$.
+Where $\text{meta}_{{m\tilde{s}}}$ indicates whether the the $m$-th progressive pattern recorded in the metadata of current instance matches the $\tilde{s}-$th element in the metadata list and $\tau$ is a learnable parameter.
+As shown in Formula (5), to reduce computational overhead we do not compress each viewpoint $v$ of the pattern representations $\{P_{nm}|n\in[1,9],m\in[1,M]\}_{v}|_{v=1}^{V}$ individually into the constructed Gaussian mixture; instead, we compute the viewpoint average $\{\bar{P}_{nm}|n\in[1,9],m\in[1,M]\}$ and plan the distribution of these averaged representations.
 Formula (5) has clearly elucidated the mechanism of SBR.
@@
 VIII-A Experiment on RAVEN
-Initially, we tested Valen on RAVEN and I-RAVEN datasets to show its potential.
-For a fair comparison, we used the same settings and equipment as RS-Tran[28] and Triple-CFN[32], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
-TABLE I: Reasoning Accuracies on RAVEN and I-RAVEN.
+We first evaluated Valen on the RAVEN and I-RAVEN datasets.
+For a fair comparison, we adopted the same settings and equipment as RS-Tran[32] and Triple-CFN[37], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
+For the RAVEN problem, the number of viewpoints $V$ in the Valen model is set to 16, corresponding to a $4\times 4$ patch segmentation.
+TABLE I: Reasoning Accuracies Achieved on RAVEN and I-RAVEN.
 Test Accuracy(%)
 Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
+CoPINet[22] 52.96/22.84 49.45/24.50 61.55/31.10 52.15/25.35 68.10/20.60 65.40/19.85 39.55/19.00 34.55/19.45
+PrAE Learner[28] 65.03/77.02 76.50/90.45 78.60/85.35 28.55/45.60 90.05/96.25 90.85/97.35 48.05/63.45 42.60/60.70
 SAVIR-T [27] 94.0/98.1 97.8/99.5 94.7/98.1 83.8/93.8 97.8/99.6 98.2/99.1 97.6/99.5 88.0/97.2
@@
 MRNet [25] 96.6/- -/- -/- -/- -/- -/- -/- -/-
+DrNet[34] 96.9/97.6 -/- -/- -/- -/- -/- -/- -/-
 RS-TRAN[32] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
@@
 Valen 99.5/99.7 100.0/100.0 99.7/99.8 98.2/99.2 99.9/100.0 99.9/100.0 99.9/99.9 98.8/99.2
+Human[22] 84.41 95.45 81.82 79.55 86.36 81.81 86.36 81.81
+GPT-4V(0-shot)[18] 12.0/- -/- -/- -/- -/- -/- -/- -/-
+GPT-4V(1-shot)[18] 12.0/- -/- -/- -/- -/- -/- -/- -/-
+Gemini Pro(0-shot)[18] 11.0/- -/- -/- -/- -/- -/- -/- -/-
+Gemini Pro(1-shot)[18] 10.0/- -/- -/- -/- -/- -/- -/- -/-
+Genome(few-shot)[19] 72.3/- 80.1/- -/- -/- 67.6/- 69.1/- -/- -/-
 The experimental records in Table I demonstrate that existing excellent RPM solvers, including Valen, exhibit varying reasoning accuracy on the RAVEN [16] and I-RAVEN [40] datasets.
@@
 Nonetheless, the same solvers still achieved different reasoning accuracies on RAVEN and I-RAVEN.
-This observation aligns with our perspective in Tine: the learning objective of probabilistic highlighting solvers is not the distribution of correct solutions, but rather the distribution delineated by both primary and auxiliary samples.
+This observation aligns with our perspective in Tine: the learning objective of probability-highlighting solvers is not the distribution of correct solutions, but rather the distribution delineated by both primary and auxiliary samples.
+In Table I, the performance of several LLMs is listed below the human benchmark [22].
+Despite high expectations, these models have failed to reach the desired threshold.
+These findings not only confirm RAVEN’s authority as a benchmark for evaluating AI capability but also demonstrate the superior performance of Valen.
 VIII-B Experiment on Bongard-Logo
-We conducted experiments on Bongard-Logo. The experimental results are presented in the table II.
-TABLE II: Reasoning Accuracies of models on Bongard-Logo.
+We conducted experiments on Bongard-Logo.
+The experiment was conducted in an environment equipped with 4 A100s graphics cards, with a batch size set to 120.
+The Adam[48] optimizer was chosen, and the initial learning rate was set to $10^{-3}$, with a per-epoch decay method applied for learning rate adjustment, featuring a decay rate of 0.999.
+It is noteworthy that the Valen model is not sensitive to these parameter settings.
+The results of the experiment are presented in Table II.
+TABLE II: Reasoning Accuracies Achieved on Bongard-Logo.
 Accuracy(%)
-Model Train FF BA CM NV
-MetaOptNet 75.9 60.3 71.6 65.9 67.5
-ANIL 69.7 56.6 59.0 59.6 61.0
-Meta-Baseline-SC 75.4 66.3 73.3 63.5 63.9
+Model Train $F$ BA CM NV
+MetaOptNet[14] 75.9 60.3 71.6 65.9 67.5
+ANIL[14] 69.7 56.6 59.0 59.6 61.0
+Meta-Baseline-SC[14] 75.4 66.3 73.3 63.5 63.9
 Meta-Baseline-MoCo[14] 81.2 65.9 72.2 63.9 64.7
-WReN-Bongard 78.7 50.1 50.9 53.8 54.3
-SBSD 83.7 75.2 91.5 71.0 74.1
-PMoC 92.0 90.6 97.7 77.3 76.0
-Triple-CFN 93.2 92.0 98.2 78.0 78.1
+WReN-Bongard[14] 78.7 50.1 50.9 53.8 54.3
+SBSD[36] 83.7 75.2 91.5 71.0 74.1
+PMoC[36] 92.0 90.6 97.7 77.3 76.0
+Triple-CFN[37] 93.2 92.0 98.2 78.0 78.1
 Valen 92.5 91.0 97.8 77.5 76.5
@@
 Funny+Valen+Tine 96.0 95.2 99.6 84.6 82.9
-Observing table II, The performance of Funny, Valen and Tine were promising.
+Human(Expert)[14] - 92.1 99.3 90.7
+Human(Amateur)[14] - 88.0 90.0 71.0
+As shown in Table II, the performance of Funny, Valen and Tine is promising.
+Notably, the transformation that adapts the Bongard Logo problem to Valen’s processing framework (illustrated in Figure 8) is more than merely an RPM-style layout strategy; it is a deliberate redesign grounded in the statistical essence of the Bongard-Logo problem.
+From a statistical standpoint, the Bongard-Logo problem involves an unknown distribution P(attribute)P({attribute}) to which the attributes of primary group images conform, whereas those of auxiliary group images do not.
+Solving the Bongard-Logo problem hinges on accurately estimating this latent distribution; the better the estimate, the higher the reasoning accuracy.
+The Tine method suggests that a probability-highlighting solver is essentially inducing a distribution that assigns high probability to primary group samples and low probability to auxiliary group samples.
+When this mechanism is applied to Bongard-Logo, its effect is equivalent to estimating P(attribute)P({attribute}), meaning that if such a solver is deployed on this task, it will inherently focus on that core statistical objective.
+Therefore, the fundamental purpose of our transformation of the Bongard-Logo problem is to externalize the core statistical task of estimating P(attribute)P({attribute}) by leveraging the intrinsic mechanism of probability-highlighting solvers, rather than merely providing a convenient interface.
+Behind the seemingly simple rearrangement of the reasoning images lies a methodology that externalizes the statistical essence of reasoning problems.
+The successful transformation of the Bongard-Logo task not only yields a new approach to the task itself but also offers guidance for other reasoning problems whose solutions hinge on estimating attribute distributions.
 VIII-C Experiment on PGM
-In this paper, we conducted experiments related to Valen on PGM. And the results are recorded in Tabel III.
-TABLE III: Reasoning Accuracies of models on PGM.
+In this paper, we conducted experiments related to Valen on PGM.
+For the PGM problem, the number of viewpoints $V$ in the Valen model is set to 16, corresponding to a $4\times 4$ patch division.
+For a fair comparison, this paper adopts the same configuration as RS-Tran[32] and Triple-CFN[37] in terms of hardware selection, data volume, optimizer parameters, data augmentation methods, batch size, and other settings.
+And the results are recorded in Table III.
+TABLE III: Reasoning Accuracies Achieved on PGM.
 Model Test Accuracy(%)
+CoPINet[22] 56.4
 SAVIR-T [27] 91.2
@@
 Valen+Tine 98.8
-Funny+Valen 99.0
+Funny+Valen 99.3
 RS-Tran+Tranclip[32] 99.0
@@
 SBR+Valen 99.4
-From the table III, it can be observed that the SBR, Tine, and Funny methods have all contributed to varying degrees of improvement in Valen’s performance on RPM problems.
-Valen combined with the Funny method is the first to achieve a reasoning accuracy exceeding 99.0% on the PGM problem without relying on any additional supervisory signals.
-Furthermore, the integration of SBR with Valen is the first approach to utilize metadata as an additional supervisory signal, achieving a reasoning accuracy of 99.4%.
-Additionally, SBR provides Valen with interpretability by endowing it with a progressive pattern extraction process.
+Human[17] >>80.0
+From Table III, it can be observed that the SBR, Tine, and Funny methods have all contributed to varying degrees of improvement in Valen’s performance on RPM problems.
+The combination of Valen and the Funny method has achieved a reasoning accuracy exceeding 99.3% on the PGM problem for the first time, without relying on any additional supervisory signals.
+Furthermore, the combination of SBR with Valen represents the first method that achieves a 99.4% reasoning accuracy by utilizing metadata as an additional supervisory signal.
 The SBR method presented in this paper is not the first to consider utilizing the accompanying metadata within RPM problem instances to assist the solver in reasoning.
-Previous works[21] have explored various approaches to leveraging metadata, but they ultimately found that forcibly imposing additional metadata learning tasks on their solvers was counterproductive.
+Previous works[25, 32, 37] have explored various approaches to leveraging metadata, but they ultimately found that directly and forcibly imposing additional metadata learning tasks on their solvers was counterproductive[25].
 RS-Tran[32] and Triple-CFN[37] adopted an indirect manner of utilizing metadata.
 Specifically, RS-Tran utilizes Tranclip[32] to pre-train its representation encoder using metadata, and Triple-CFN required a similar warm-start process.
-However, the SBR method in this paper is the first to directly utilize metadata to model the distribution of progressive pattern representations, which is an unexplored approach in previous works and has achieved promising results.
-We also conducted experiments on the interpolation and extrapolation problems in PGM, and the results are recorded in Table IV.
-As can be seen from the table IV, the Funny method brings a remarkable improvement to Valen in interpolation.
-TABLE IV: Generalization Results of Valen in PGM.
-Accuracy(%)
-Dataset Valen Valen+Tine Funny+Valen SBR+Valen
-Interpolation 81.2 82.8 87.6 92.5
-Extrapolation 18.4 18.5 18.5 12.8
+However, the SBR method in this paper is the first to directly utilize metadata to model the distribution of progressive pattern representations, which is an unexplored approach in previous works and has achieved remarkable results.
+We also carried out experiments on the generalization issue of PGM, and the findings are documented in Table IV.
+As evident from Table IV, the Funny method significantly enhances Valen’s performance on the attribute generalization task Interpolation, whereas SBR exhibits notable advancements in the two progressive generalization patterns, Held-out Pairs of Triples and Held-out Attribute Pairs.
+TABLE IV: Reasoning Accuracies Achieved on PGM Generalization Tasks.
+Dataset
+Model/
+Task
+Interpolation Extrapolation
+Held-out
+Attribute
+shape-color
+Held-out
+Attribute
+line-type
+Held-out
+Triples
+Held-out
+Pairs of
+Triples
+Held-out
+Attribute
+Pairs
+MRNet[25]
+68.1 19.2 16.9 30.1 25.9 55.3 38.4
+RS-Tran[32]
+77.2 19.2 12.9 24.7 22.2 43.6 28.4
+Triple-CFN+Re-space layer[37]
+80.4 18.4 12.6 25.2 22.0 44.5 29.2
+Valen 81.2 18.4 12.8 26.0 22.8 44.8 29.5
+Valen+Tine
+82.8 18.5 13.2 26.4 23.3 45.2 30.0
+Funny+Valen
+87.6 18.5 13.6 27.2 24.1 45.2 30.2
+SBR+Valen
+92.5 12.8 13.8 26.8 28.2 98.1 98.2
+Table IV also indicates that MRNet demonstrates a slight advantage in the remaining generalization tasks; however, this paper takes a cautious stance, arguing that it is difficult to unequivocally determine which solver is better when the reasoning accuracy hovers around 30%, as a reasoning accuracy below this threshold already underscores the limitations of the solvers, making it inadequate to discuss superiority at this point.
 IX Conclusion
@@
 To bridge the discrepancies between the estimated solution distribution and the true correct solution distribution, the paper introduces three methods: Tine, Funny, and SBR.
-The Tine method, based on adversarial learning, aims to assist Valen in estimating a solution distribution closer to the correct one.
-The Funny method models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution.
-The SBR method further extends this idea by modeling the distribution of progressive patterns representation as a mixture of Gaussian distributions.
-This paper demonstrates the importance of explicitly planning the distribution of solutions in solving visual abstract reasoning problems through the successful application of the SBR, Funny, Valen and Tine.
+  1. 1.
+    The Tine algorithm conducts an in-depth analysis of the reasoning mechanism of probability-highlighting solvers and elucidates that the inadequate quality of auxiliary samples in reasoning problem instances often leads to the estimated distribution deviating from the correct solution distribution when using such end-to-end deep learning models to estimate the solution distribution for abstract reasoning problems.
+Consequently, the Tine method ingeniously incorporates the concept of adversarial learning into the solver’s training process, leveraging adversarial techniques to generate more judicious auxiliary samples, for the Valen model, thereby refining its solution distribution estimation process.
+Tine’s contributions lie not only in demonstrating the effectiveness and potential of adversarial learning frameworks in solver training, but also in revealing the crucial role of auxiliary samples in delineating the boundaries of probability-highlighted solver solution distributions, and highlighting the importance of the quality of negative examples for machine intelligence in learning abstract concepts.
+  2. 2.
+    The Funny method, embracing the spirit of the Tine algorithm, introduces additional reasoning image regression tasks for Valen to achieve mutual information supervision, thereby modeling the distribution of reasoning sample representations encoded by Valen as a Gaussian mixture distribution to enable it to more directly capture the true form of the correct solution distribution.
+The Half-Split Decoder designed in the Funny method has also made significant contributions to balancing regression and reasoning tasks, potentially opening up new opportunities for solving reasoning problems in the form of Image Generation Model.
+The Funny method emphasizes the importance of modeling the representation distribution of reasoning samples for enhancing the level of solver’s reasoning intelligence.
+  3. 3.
+    The SBR method further extends the idea of the Funny method.
+By modeling the distribution of Valen’s encoded progressive pattern representations as a more interpretable and orderly mixture of Gaussian distributions, the SBR method significantly enhances Valen’s reasoning accuracy and interpretability.
+The SBR method represents a novel attempt in the field of RPM solver training, embodying an innovative approach and mechanism to utilize metadata.
+It successfully reverses the previously observed phenomenon where directly introducing metadata as a supervisory signal to the solver actually impairs the reasoning process[25, 32, 37].
+The SBR method emphasizes that for solvers dedicated to learning abstract concepts, it is crucial to meticulously plan the distribution of latent variables aimed at extracting these abstract concepts, as this significantly enhances their reasoning intelligence.
+This paper demonstrates the importance of explicitly planning the distribution of solutions in solving visual abstract reasoning problems.